### v1

# Towards Better Comprehension of Breaking Changes in the NPM Ecosystem

Abstract.
Code evolution is prevalent in software ecosystems, which can provide many benefits, such as new features, bug fixes, security patches, etc., while still introducing breaking changes (BCs) that make downstream projects fail to work. Breaking changes cause a lot of effort to both downstream and upstream developers: downstream developers need to adapt to breaking changes and upstream developers are responsible for identifying and documenting them. In the NPM ecosystem, characterized by frequent code changes and a high tolerance for making breaking changes, the effort is larger.
For better comprehension of breaking changes in the NPM ecosystem and to enhance breaking change detection tools, we conduct a large-scale empirical study to investigate breaking changes in the NPM ecosystem. We construct a dataset of explicitly documented breaking changes from 381 popular NPM projects. We find that 93.6% of the detected breaking changes can be covered by developers’ documentation, and about 19% of the breaking changes cannot be detected by regression testing. Then in the process of investigating source code of our collected breaking changes, we yield a taxonomy of JavaScript and TypeScript-specific syntactic breaking changes and a taxonomy of major types of behavioral breaking changes. Additionally, we investigate the reasons why developers make breaking changes in NPM and find three major reasons, i.e., to reduce code redundancy, to improve identifier name, and to improve API design, and each category contains several sub-items.
We provide actionable implications for future research, e.g., automatic naming and renaming techniques should be applied in JavaScript projects to improve identifier names, future research can try to detect more types of behavioral breaking changes. By presenting the implications, we also discuss the weakness of automatic renaming and BC detection approaches, such as the lack of support for public identifiers and various types of breaking changes.
Breaking Change, NPM, JavaScript, Code Evolution
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Software evolution
1. Introduction
The evolution of code is prevalent in software ecosystems (Konstantopoulos et al., 2009; Raemaekers et al., 2012; Lehman, 1980; Traini et al., 2021). Developers of upstream software libraries make code changes to incorporate new features, bug fixes, security patches, component refactorings, and extra-functional improvements (Traini et al., 2021; Tao et al., 2012). However, code evolution may break the contract previously established with its downstream by introducing breaking changes (BCs) in its public APIs, making client applications fail to work (Xavier et al., 2017). For example, renaming frequently used methods or classes can make client projects that depend on the code fail to compile. Therefore, downstream developers will be required to make efforts to adapt to such BCs.
Considering the risks and the effort brought by BCs to downstream developers, many works have been utilized to help developers analyze and measure the impacts of BCs (Mezzetti et al., 2018; Møller and Torp, 2019; Don, [n. d.]). However, they relied on test suites in downstream projects, and either detected JavaScript BCs by directly running test suites (Don, [n. d.]), or via API models [^1] Intuitively, an API model defines the type restrictions of an API (Mezzetti et al., 2018). generated from dynamic execution of test cases (Mezzetti et al., 2018; Møller and Torp, 2019). This leads to the result that BCs observed in prior studies are located in well-known APIs that are popularly used in downstream projects, such as “Client.socket” in socket.io [^2] https://github.com/socketio/socket.io/commit/b73d9be, “Request.prototype.onResponse” in request [^3] https://github.com/request/request/commit/d05b6ba, and each in async. However, BCs are reported to upstream projects because they already manifested themselves in downstream test cases or already caused bugs. For example, in version 3.0.0 of a famous JavaScript project Lodash, the behavior of function “mixin” was changed and not explicitly marked by upstream developers, and a downstream developer reported this issue by self-constructed test code. [^4] https://github.com/lodash/lodash/issues/880 Hence, completely relying on test case execution is not sufficient for understanding BCs.
We observe that in the NPM ecosystem, a number of projects write commit messages complying with Conventional Commits (Con, [n. d.]). According to Conventional Commits, “BREAKING CHANGE” tokens are documented in commit messages to indicate that the corresponding commits contain breaking changes. This is because the core developers of upstream libraries need to ensure the stability of the API and protect the reputation of the upstream projects. When a new code change is submitted for review, they carefully inspect it for the presence of BCs, as well as other bugs or defects. By leveraging the documented BCs provided by developers, there is a potential opportunity to notify downstream users about these BCs.
However, developers experience increased pressure within the NPM ecosystem, primarily driven by frequent code changes and a high tolerance for BCs (Bogart et al., 2021, 2016). Another possible factor is that different developers may have different perceptions of BCs and some of them may ignore the BCs, leading to unaware BCs and causing unforeseen bugs in downstream projects. If we could characterize the large number of documented BCs at source code level, it would assist upstream developers in gaining a deeper understanding of the types of commits that could potentially introduce BCs. Furthermore, such categorization would provide valuable insights for researchers to develop tools aimed at detecting BCs in the future.
To bridge this gap, we conduct an empirical study to better comprehend documented BCs. We first select 381 popular JavaScript projects (randomly selected from 35,786 projects with more than 50 GitHub stars in the NPM ecosystem) and clone their associated repositories. We then identify BC-related commits admitted by developers by searching for BREAKING CHANGE tokens in commit messages. For the 5,242 identified BC commits, we manually select 1,519 commits that 1) containing not too long commit messages, 2) related to JavaScript production code and 3) associated with documentation that can explain the reasons behind the commit. These selected BC commits are distributed across 131 distinct projects and the projects can be grouped into six categories according to functionalities and usages, i.e., utilities, frontend projects, development-related tools, database-related tools, plugins and Web API related tools (detailed in Section 3), which makes the BC commits in our projects representative. To extract BCs from these BC commits, we build BC type categories by learning from a previous API detection tool for Java (Brito et al., 2018) and adding new JavaScript and TypeScript-specific features. Our taxonomy includes Remove, Rename, Change Signature, Change Behavior types, etc., which is detailed in Section 2.2. Through several thematic analyses, we obtain JavaScript and TypeScript specific code features (RQ2 and RQ3) and a taxonomy of reasons behind BCs. We answer the following research questions in this study:
RQ1: To what extent do detected breaking changes and documented breaking changes overlap? In our work, we collect BCs from developers’ intention, i.e., the explicitly documented BCs. We first check whether documented BCs actually break test code. Therefore, we do regression testing on our collected BC commits, similar to previous works (Mezzetti et al., 2018; Møller and Torp, 2019; Mostafa et al., 2017; Zhang et al., 2022). We find 97% of detected BCs are documented and 81% of the documented BCs can be detected by regression testing. The result shows that most detected BCs are well documented, while a proportion of documented BCs cannot be detected by regression testing. Since the documented BCs cover most of the detected BCs, hence extracting BCs from documentation (especially commit messages, issues, and pull requests) is reasonable.
RQ2: What syntactic breaking changes in the NPM ecosystem are specific to JavaScript and TypeScript? We investigate whether there are some BCs related to JavaScript-specific language features. Despite the triviality of most syntactic BCs (such as moving classes to another package, removing a field of a class), we find some JavaScript-specific BCs are notable, including two removing operations, i.e., 1) removing default export, 2) removing export of an element in one place, and five types of JavaScript and TypeScript specific signature changes, e.g., parameter changes in configuration objects, switch between callback and Promise, etc.
RQ3: How do developers make behavioral breaking changes? In this RQ, we investigate how developers perform BCs at source code level. Since syntactic BCs can be detected through syntax analysis and code refactoring detection, we mainly focus on behavioral BCs (a.k.a, semantic BCs in some works) that change the internal program logic. We find four major types of Change Behavior BC, i.e., 1) changing the specification return values, 2) changing process for some option values, 3) changing default or initial value of variables and 4) changing error handling method.
RQ4: Why do developers make breaking changes in NPM ecosystem? In this RQ, we revisit the rationale of BCs in the NPM ecosystem and find three main factors (each with some sub-factors) that motivate developers to make BCs, i.e., 1) to reduce code redundancy, 2) to improve identifier name, 3) to improve API design, and divide each reason into some sub-items, which extended the previous works (Bogart et al., 2016, 2021; Brito et al., 2020).
Based on our empirical findings, we provide actionable implications for future research, including 1) automatic renaming deserves much concern in JavaScript projects since poor identifier name is a main contributor to technical debt, 2) automatic tools for ensuring code consistency are necessary, 3) future research can try to detect more types of behavioral BCs. By presenting these implications, we also discuss the weakness of automatic renaming and BC detection approaches, such as lacking support for public identifiers and various types of BCs.
The contributions of our paper are two-fold:
  1. (1)
    We build a carefully constructed breaking change dataset extracted from a wide range of JavaScript and TypeScript projects in the NPM ecosystem.
  2. (2)
    We empirically identify how developers perform BCs at source code level, yielding notable findings like JavaScript and TypeScript specific syntactic BCs and typical actions of behavioral BCs. We also investigate why developers perform BCs in the NPM ecosystem.
The remainder of this paper is organized as follows. Section 2 introduces preliminary knowledge and motivating examples. Section 3 presents the methodology of our study. Section 4 details our empirical findings. Section 5 provides our implications for future research from this study. Section 6 discusses the threats to validity. Section 7 reviews the related work and Section 8 concludes our work.
Data availability. We provide the datasets of our research at https://zenodo.org/records/13219136.
2. Background
In this section, we describe some preliminary knowledge on breaking changes.
2.1. Preliminary of Breaking Changes
Prior to our work, a number of studies have uncovered BCs from many aspects, such as the occurrence of BCs, the impacts of BCs on downstream projects, and motivations for making BCs. The research works (Decan and Mens, 2019; Venturini et al., 2023; Raemaekers et al., 2017) have demonstrated the widespread occurrence of BCs in the NPM ecosystem and their effects on client applications. Venturini et al. found that in their sampled packages, 11.7% of all client packages and 13.9% of their releases are impacted by breaking changes, and notably, 44% of the breaking changes are introduced in minor or patch releases (according to Semantic Versioning (Sem, [n. d.]), developers should not introduce breaking changes in non-major releases) (Venturini et al., 2023). Bogart et al. found that coarse-grained motivations for making BCs include requirements and context changes, bugs and new features, rippling effects from upstream changes, and technical debt from postponed changes (Bogart et al., 2016, 2021).
However, despite much knowledge of BC, to the best of our knowledge, none of the studies precisely defined BC. Researchers collected BCs in various forms in previous studies. For example, in the evaluation of several breaking change detection tools (Zhang et al., 2022; Mezzetti et al., 2018), researchers collected BCs by running downstream test cases: if a test case of a downstream project could not pass with a newer provider (after the code change), then the code change was identified as a BC. However, on the one hand, not all providers are dependent on many client applications. If a code change has no triggering test case in client applications, we cannot determine whether it is actually incompatible. On the other hand, client code may not follow the specification of providers and incorrectly access APIs, e.g., passing an improper value to a function, which will result in undefined behavior (Møller and Torp, 2019), including test failure, but it does not reflect a BC. In another breaking change detection tool APIDiff (Brito et al., 2018), Brito et al. utilized code refactoring detection tool RefDiff (Silva and Valente, 2017) to identify syntax-related BCs (e.g., changing method signature or renaming class), however, as they pointed out, the detected BCs are just breaking change candidates (BCC), since some syntax changes are applied to internal methods and classes, which are not intended for public use. To this end, they also asked developers to check whether the detected BCCs are actually BCs (Brito et al., 2020).
In our study, we adopt Brito et al.’s criteria (Brito et al., 2020), since we also investigate the BCs from developers’ perspective, i.e., how and why they perform BCs. Specifically, a BC should 1) be confirmed by developers, 2) be categorized into pre-defined BC types. To achieve the first criterion, we identify BCs by searching explicit BC declarations in the documentation (usually commit messages). To achieve the second criterion, we define several BC types for JavaScript and TypeScript on the basis of Brito et al.’s Java BC types (Brito et al., 2018) and JavaScript development experience. We detail the BC types in Section 2.2 and describe how to extract BCs in Section 3.3.
2.2. Types of Breaking Changes in NPM Projects
Table 1. Syntactic Breaking Change Types Adopted from Previous Works
Type Supported Elements
Rename module, class, interface, enum, type, method, field, constant, variable
Remove module, class, interface, enum, type, method, field, constant, variable
Move module, class, interface, enum, type, method, field, constant, variable
Inline method
Push Down method, field
Change Signature method, field
Existing BC detection tools for Java, typically APIDiff (Brito et al., 2018), support many syntactic and object-oriented programming (OOP) related BC types, such as remove classes and push down fields. Since OOP features are also supported in JavaScript (since ECMAScript 2015 (ECM, [n. d.]), class can be directly used while legacy code must leverage prototype mechanism (inh, [n. d.])), we learn from the BC types for Java language and make them adapt to JavaScript and TypeScript (a superset of JavaScript, in practice many source code files are written in TypeScript and will be compiled into JavaScript files) by retaining available BC actions in JavaScript and adding more BCs which are not considered previously. We then construct a category of JavaScript and TypeScript BCs (shown in Table 1), where interface, enum and type are only available in TypeScript. We explain each type as follows:
  1. (1)
    Rename BC refers to identifier changes of public code elements. Rename BCs can be applied to almost all code elements, such as exported classes, interfaces, and enums, as well as methods and fields in exported classes and interfaces.
  2. (2)
    Remove BC refers to removal of public code elements. Remove BC can be applied to almost all code elements. After a Remove BC, the affected public element is not accessible by downstream applications, e.g., an exported class is completely removed, or no longer marked as exported.
  3. (3)
    Move BC can be applied to all APIs supported by Rename BCs. For example, a class is moved to another package since its functionality is more related to that package.
  4. (4)
    Inline is a refactoring operation in object-oriented programming: removing a public method and copying its body into an existing method. Simple methods are often inlined to make code more straightforward. In fact, Inline can be regarded as a special case of Remove, since the inlined public method is removed from public access. By contrast, normal “Remove” actions do not copy the content of the removed method to another existing method. However, since inline is a very common refactoring operation in Object-oriented languages, we still regard it as a separate type of BC, like previous works (Brito et al., 2018; Silva and Valente, 2017).
  5. (5)
    Push Down is also a refactoring operation in object-oriented programming: moving the method in the parent class into the child class since this method is only used in one child class in reality. Push down can be regarded as a special case of Remove BC.
  6. (6)
    Change Signature BC refers to the modification of method signatures except for directly changing method names. Adding or removing modifiers (such as “static”, “private”) and changing the parameters are typical refactoring actions of this type.
Besides syntactic BCs, Change Behavior BC (also called semantic BC in some works (Mezzetti et al., 2018; Zhang et al., 2022)) refers to those operations changing the internal behavior (e.g., programmatic logic in methods, and content of global constants that can be used by some methods) rather than syntactic elements, such as changing class names and adding required parameters. In other words, after a behavioral BC, the affected classes (or methods, interfaces, etc.) can be invoked as the same way as before.
In our work, we focus on BCs in JavaScript source code, hence we do not take BCs in non-source code files into consideration, typically CSS, Markdown, HTML, package.json files and configuration files related to TypeScript, Webpack and ESLint, etc.
3. Methodology
3.1. Data Collection
Due to the vast number of JavaScript projects in NPM, our study focuses on selecting the most popular projects to ensure that the collected BCs are representative. To do so, we utilize the Libraries.io open source repository and dependency metadata provided by Reid et al. (Reid, 2020) to retrieve the most popular JavaScript and TypeScript projects since Libraries.io is commonly used in prior studies (He et al., 2021; Decan et al., 2019, 2018). We first select popular projects having more than 50 GitHub stars and associated with an existent repository (deprecated and unmaintained repositories are not considered), which yields 35,786 repositories. We then randomly select 381 repositories (95% confidence level and 5% margin of error) for further investigation.
Since we consider the BCs that are explicitly confirmed by developers, we try to extract BCs from developers’ documentation, typically commit messages, changelogs, issues and pull request text on GitHub. We notice that Conventional Commits (Con, [n. d.]) provides a standard for writing readable commit messages. According to the standards, if a commit message contains a BREAKING CHANGE section, the commit is identified as a breaking change. For example, the text below shows a typical commit message indicating a breaking change:
refactor: compiler -> runtimeCompiler
BREAKING CHANGE: compiler option has been renamed to runtimeCompiler
We first check whether Conventional Commits are widely used, analyzing the commits from the 381 projects we used and find that We find that 360 out of 381 repositories contain commits that conform to Conventional Commits and 195 projects have over 80% commits that follow Conventional Commits. We use the regular expression below to check Conventional Commit compliance:
[…]
This indicates that Conventional Commit is widely used and we can use it as a simple way to extract plenty of BCs. Considering that many projects do not follow this standard, we also try our best to extract BC commits from changelogs of each project (if developers have declared), issues and pull requests on GitHub. For example, the project tj/commander.js does not use Conventional Commits, then we consult the mentioned pull request IDs (and commits related to the PR) in CHANGELOG.md file [^5] https://github.com/tj/commander.js/blob/master/CHANGELOG.md, shown in Figure 1. In this way, we obtain 5,242 commits.
Figure 1. Screenshot of the Changelog of tj/commander.js (Breaking Changes Are Highlighted)
Then two of the authors manually inspect them and remove some of them that are not related to JavaScript production code (e.g., only modifying package.json, Markdown documentation, build scripts, test cases) or contain very long commit messages (over 10 lines) that are difficult to understand. For example, if a commit only states that “drop support for Node.js 14” and modifies the corresponding item in package.json, it will be discarded. Additionally, we do not include the commits that contain multiple breaking change declarations. Specifically, we remove the 965 commits that only updated dependency versions, 456 commits that only changed minimum versions of supported platforms, and 244 commits that only changed build scripts or JSON configuration files. We also remove 865 commits that contain over 10 line commit messages or multiple BREAKING CHANGE declarations. Finally, we retain 2,724 candidate commits.
3.2. Regression Testing
We use regression testing to detect BCs in our selected projects (381 in total) to check:
  1. (1)
    what percentage of detected BCs are well documented,
  2. (2)
    what percentage of documented BCs can be detected by regression testing.
To check ①, due to the number of the total commits is too large (1,455,003), we first sample 16,453 commits for further analysis (99% confidence level and 1% margin of error). We then remove the commits in these projects that only contain non-JavaScript code changes (e.g., test code changes, Markdown and HTML documentation changes, dependency updates in package.json). Then for each commit $c$ and its prior commit $c^{\prime}$ we do regression testing following the steps below:
  1. (1)
    We run the test cases in $c^{\prime}$ directly (using npm test command). We discard the commits that cannot be configured (e.g., there are version conflicts during installation).
  2. (2)
    We restore the production code from $c$. The production code refers to JavaScript and TypeScript source code files not located in test directory.
  3. (3)
    We run the test cases (using npm test command).
  4. (4)
    If the first step succeeds and the third step fails, then we say that $c$ is a detected BC commit.
  5. (5)
    If a detected BC commit is documented with BREAKING CHANGE, or can be found in changelogs, issues, or pull requests, we regard the commit as a documented BC commit.
We use the test cases written by developers rather than dependents since 1) we have found that all of the 381 projects provide test suites and developers’ test cases can better express the intended usage of APIs (Dinella et al., 2022), while many of the 381 projects do not have plenty of dependents, 2) some projects are not intended for pragmatic use, especially CLI tools in our projects such as npm/cli, therefore no clients access the APIs in these projects.
To check ②, we apply the steps described above to the 2,724 BC commits from Section 3.1.
3.3. Breaking Change Selection and Analysis
3.3.1. Breaking Change Selection and Categorization
We select feasible BCs for further research from the 2,724 candidate BC commits and categorize them into types described in Section 2.2. Since we need to understand what a commit changed and why developers made it, we require the commit message and other documentation of a commit (including issues, pull requests and changelogs) to contain some explanation of why developers made that commit. We here reuse Tian et al.’s taxonomy of why information in commit messages (Tian et al., 2022). Typical why information include:
  1. (1)
    Issue description: about the linked issue, weakness of current code implications, etc.
  2. (2)
    Requirements Illustration: about the usage need, out-of-date statement, etc.
  3. (3)
    Objective description: objectives, such as fixing bugs, improving performance, refactoring code, etc.
  4. (4)
    Necessity Implication: relation to prior commits, benefits of making such code changes, etc.
For example, considering the commit cebd670a of angular/angular [^6] https://github.com/angular/angular/commit/cebd670a, the developers just stated renaming the method requestCheck in ChangeDetectorRef to markForCheck. By inspecting the related issue (#3403), the developer said that “When I first saw the requestCheck() method on ChangeDetectorRef, I assumed it was how I manually run change detection. Others have assumed this as well”, which is a description of “Necessity Implication”.
If there is no why information covering the aspects mentioned above, we discard the commit. Two of the authors independently check whether the documentation contains reason-related information (Cohen’s Kappa value is 0.80) and determine the type of each commit (using the taxonomy in Section 2.2, Cohen’s Kappa value is 0.77). Then the authors hold meetings to solve the disagreements. In total we identify 1,519 commits that contain reason-related information.
3.3.2. Breaking Change Distribution
Figure 2. Categories of Our Investigated Breaking Changes
Figure 2 presents the percentages of different types of BCs. It can be obviously seen that Change Behavior BCs make up the largest proportion of all BC types. By contrast, Move BCs make up the smallest percentage. In our further investigation, we mainly focus on Change Behavior BCs.
Table 2. NPM Projects Containing Breaking changes Used in Our Study
Category Description Number
Frontend Used in Web browser environment 45
Web API Providing access to RESTful APIs 9
Database Tools Providing database operations 4
Development Tools Providing project management 25
Plugin Served as plugins for other NPM projects 3
Utility Other useful JavaScript libraries 45
Table 3. Proportion of Involved Projects of Each Breaking Change Category
BC Category Frontend Utility Development Database Web API Plugin # All
Remove 10/45 12/45 8/25 4/4 4/9 0/3 38/131
Rename 10/45 10/45 3/25 3/4 3/9 0/3 29/131
Move 2/45 2/45 1/25 1/4 1/9 0/3 7/131
Inline 0/45 0/45 1/45 0/4 0/9 0/3 1/131
Change Signature 8/45 10/45 6/25 3/4 2/9 0/3 29/131
Change Behavior 44/45 41/45 25/25 4/4 9/9 3/3 130/131
The projects that contain BCs in our study serve various types of functionalities, including command-line interface utilities, REST API SDKs and Web frontend frameworks. They can be grouped into six categories (shown in Table 2). For example, pnpm/pnpm [^7] https://github.com/pnpm/pnpm is a alternative CLI tool for official NPM implementation, renovatebot/renovate provides automatic project build and aws/aws-cdk [^8] https://github.com/aws/aws-cdk wraps the AWS cloud APIs, gajus/eslint-plugin-flowtype [^9] https://github.com/gajus/eslint-plugin-flowtype is a ESLint plugin, and async/async [^10] https://github.com/async/async is a fundamental utility of asynchronous programming. Table 3 shows the number of involved projects of each type of BCs.
3.3.3. Breaking Change Labeling
For the 1,519 remaining commits, two of the authors conduct two thematic analyses to analyze ① source code features (for RQ2 and RQ3) and ② reasons behind the commit (for RQ4). We use the guideline recommended by Cruzes et al. (Cruzes and Dyba, 2011). The guideline for the two analyses is shown below. Two of the authors independently perform the steps above, and they hold a series of meetings to solve the agreements.
  1. (1)
    We read the related documentation, especially commit messages carefully to understand what the developers want to express.
  2. (2)
    We read the documentation again and generate phrases as initial codes that describe the source code features, how developers make breaking changes and the reasons behind the commit.
  3. (3)
    We aggregate the codes with similar meanings and generate a theme name to describe each cluster.
  4. (4)
    We then review all themes and try to merge the themes with similar semantics, or make the similar themes become sub-items of a new theme.
After we obtain all themes, we assign possible themes to each BC. Note that for simplicity, we assign each retained BC with one reason, which is consistent with the previous works on motivations behind breaking changes (Brito et al., 2020). In this process, we held several meetings to resolve the disagreements on the reason behind each commit since some BC commits might have multiple potential reasons. For example, in the commit 2713380 of reactivex/rxjs [^11] https://github.com/reactivex/rxjs/commit/2713380, developers renamed inspect to audit, the reasons behind this rename operation was presented in Issue #1505121212https://github.com/reactivex/rxjs/issues/1505 and #1387131313https://github.com/reactivex/rxjs/issues/1387. Developers mentioned two possible reasons, i.e., name collision (since inspect conflicts with util.inspect provided by Node.js [^14] See explanation in Issue 1387 of Rxjs.) and synonym replacement (since audit is synonymous to inspect and it is less confusing). By carefully reading and understanding the related issues, two of the authors reach an agreement and assign the reason “to avoid name conflict” to this BC, because if inspect were not used in Node.js native objects, the identifier inspect is still acceptable and does not need to be renamed. In our dataset, the authors have very few disagreements due to multiple potential reasons.
4. Results
In this section, we present our empirical results for each research question.
4.1. RQ1: To What Extent do Detected Breaking Changes and Documented Breaking Changes Overlap?
From 16,453 commits used in regression testing in Section 3.2, we detect 295 BC commits (1.8%), and 276 of them (93.6%) are documented by developers. And for 2,724 BC commits, 2,206 (81%) can be detected by regression testing.
The results illustrate that most detected BCs are well documented, while a proportion (19%) of documented BCs cannot be detected via regression testing since developers might forget to write test cases or current test cases are too general to cover the code modifications.
The undocumented BC commits are possibly due to ① the commit is made too long ago when Conventional Commits was not proposed, and ② the project does not comply to Conventional Commits. And the main reasons that regression testing cannot detect BCs is that the BC behavior can only be triggered by external conditions (e.g., network error), hence it is difficult to simulate such a situation. For example, in commit 61e7a81a of octokit/rest.js, the “followRedirects” option is no longer supported. This BC can only manifest itself when the HTTP server returns a status code between 301 and 307. However, this is not often seen, and developers have not written any test case to cover the process of “followRedirects” option. Developers might forget to write test cases and current test cases might be too general to cover the code modifications, hence automatic test case generation may be useful.
For the 19% of BCs that cannot be detected, since developers in popular projects have rich development experience, we believe those BCs are trustable and worth attention. Although some BCs currently do not affect many downstream projects, they should still be pointed out since in some cases downstream developers may write code that invokes the broken API, then the downstream developers can confirm the breaking changes quickly by looking up commit messages. For example, in commit 669592d of Socket.io [^15] https://github.com/socketio/socket.io/commit/669592d contains a breaking change that removed Socket#binary method. It is not detected when we run regression testing. However, after several months, one developer asked for an alternative method of Socket#binary [^16] https://github.com/socketio/socket.io/discussions/3826. This indicates that although an API in a BC is not frequently used, and currently no test cases can trigger it, the API is possibly invoked at a certain time.
4.2. RQ2: What Syntactic Breaking Changes in NPM Ecosystem Are Specific to JavaScript and TypeScript?
While most code-level actions in BCs are also available in other programming languages, such as changing parameter orders and removing classes, etc., we highlight the noticeable syntactic BC actions in our investigated projects.
4.2.1. JavaScript-specific remove operations
In ECMAScript 6, a module can have many exported items and a default export item. Therefore, developers can not only remove the code of a class, interface or enum, etc. like other OOP languages, but also just remove them from exported list and put them into internal source code files. We present the JavaScript-specific remove operations in our collected BC commits as follows:
  1. (1)
    Remove a default export (6/211 cases). For example, in commit 102e4b0 of nodkz/mongodb-memory-server, developers simply removed default export of the module util. Therefore, the code import generateDbName from 'util' will not work.
  2. (2)
    Remove an export position (17/211 cases). For example in commit cfbfaac of reactivex/rxjs, the types in rxjs/interfaces modules are no longer accessible and users must use them by importing them from the main module. Therefore, users can import UnaryFunction with the following two methods, while after the code change, they can only import UnaryFunction using the second form.
    import { UnaryFunction } from ’rxjs/interfaces’;
    import { UnaryFunction } from ’rxjs’;
4.2.2. JavaScript-specific signature changes
Most signature changes in JavaScript are also available in other languages like Java, e.g., adding, removing, and reordering parameters. However, some JavaScript-specific signature changes should be focused on. We present the categories as follows:
  1. (1)
    About parameters in configuration objects (15/134 cases). In JavaScript, many properties can be packaged in one configuration object. Considering the diff shown below, the second parameter of the method “userinfo” is an object, and this function only extracts four properties in this object using destructing assignment syntax (des, [n. d.]). After the code change, the “verb” property is renamed to “method”, since “GET” is an HTTP method. Correspondingly, the internal code related to this property is also changed.
    async userinfo(accessToken, {
    - verb=’GET’, via=’header’, tokenType, params,
    + method=’GET’, via=’header’, tokenType, params,
  2. (2)
    About this parameter (6/134 cases). In the old implementation, the method accepts a callable object (typically a function) and this, and binds the function to this171717bind is a native function in JavaScript. See https://developer.mozilla.org/en-US/docs/Web/JavaScript/Reference/Global_Objects/Function/bind. After the code change, the method no longer provide this parameter, the user should first bind the callable object to this. For example, in the code snippet shown below, findIndex only accepted predicate parameter.
    export function findIndex<T>(
    predicate: (value: T, index: number, source: Observable<T>) => boolean,
    - thisArg?: any
    ): OperatorFunction<T, number> {
    - return operate(createFind(predicate, thisArg, ’index’));
    + return operate(createFind(predicate, undefined, ’index’));
  3. (3)
    About sync and async (20/134 cases). In JavaScript, asynchronous functions can be implemented by using callback arguments and Promise objects. Legacy JavaScript asynchronous functions are designed with callback-style interface, while new implementations often use async directly. For example, in the commit 8c3cecae of Mongoose, developers provided multiple signatures of an API before the code change, e.g., createIndex have the following function signature:
    1createIndexes(options: mongodb.CreateIndexesOptions, callback: CallbackWithoutResult): void;
    2createIndexes(callback: CallbackWithoutResult): void;
    3createIndexes(options?: mongodb.CreateIndexesOptions): Promise<void>;
    After the code change in this commit, the signature 1 and 2 were removed. The code diff below shows this pattern:
    - SomeFunction(param, (returnValue, err) => {
    - // Some code about returnValue
    + let returnValue = await SomeFunction(param);
    + // Some code about returnValue
    Another case is converting synchronous to asynchronous (or vice versa), e.g., simply adding async modifier to the function signature.
  4. (4)
    About undefined and null (8/134 cases). The keyword null represents an intentional empty value, while undefined occurs in accessing uninitialized variables or non-existent object properties, and should not be explicitly assigned to an object. The diff below shows this pattern:
    - public get(key: Key): Value | null {
    - return this.collection.get(key) || null;
    + public get(key: Key): Value | undefined {
    + return this.collection.get(key);
    However, in JavaScript projects without d.ts files, there are also signature related changes, e.g., a method no longer accepts a null value, or a null value might cause error after a code change. In this case, developers cannot represent the change in method signature, therefore, we classify the case as behavioral change.
  5. (5)
    Switch between “required” and “optional” of fields in method or interface declarations (3/134 cases). TypeScript script projects and part of JavaScript projects use d.ts files to declare interface and method signature. For example, if one parameter is changed to be required, then a “?” will be removed from the parameter type. The diff below (from commit 86074a6 of coinbase/rest-hooks) shows the signature change type:
    static url<T extends typeof SimpleResource>(
    this: T,
    - urlParams?: Partial<AbstractInstanceType<T>>,
    + urlParams: Partial<AbstractInstanceType<T>>,
    ): string {
    - if (urlParams) {
4.3. RQ3: How do Developers Make Behavioral Breaking Change?
We identify four main types of behavioral changes in our collected JavaScript projects of NPM ecosystem. Each type is presented as follows:
  1. (1)
    Changing the specification of return values (79/1,034 cases). One typical case is that the type of returned objects are changed. For example, in commit b99f6d3 of automattic/mongoose, MongooseArray.map() returned a plain JavaScript array rather than a headless Mongoose array:
    map() {
    - const ret = super.map.apply(this, arguments);
    - ret[arraySchemaSymbol] = null;
    - ret[arrayPathSymbol] = null;
    - ret[arrayParentSymbol] = null;
    + const copy = [].concat(this);
    - return ret;
    + return Array.prototype.map.apply(copy, arguments);
  2. (2)
    Changing the process of some options (231/1,034 cases). Suppose that one possible value of an option (can be in JSON configuration or parameters) is not supported, then the code that gets properties from the option will be changed, and the corresponding branch of a specific option value will be removed. For example, in commit ad4f1493 of project octokit/rest.js [^18] https://github.com/octokit/rest.js/commit/ad4f1493 (shown below), the judgment of options.type is changed, and netrc branch in a switch structure is also removed to completely remove support for netrc mechanism.
    if (!options.type ||
    - ’basic|oauth|client|token|integration|netrc’.indexOf(options.type) $=$ -1
    + ’basic|oauth|client|token|integration’.indexOf(options.type) $=$ -1
    throw new Error("Invalid authentication type, must be" +
    - ’basic’, ’integration’, ’oauth’, ’client’ or ’netrc’"
    + ’basic’, ’integration’, ’oauth’, or ’client’"
    In CLI tool projects, developers often change the process of some options to remove or change the purpose of an option. For example, in commit f6fd0c3 of pnpm/pnpm, the option --store-path is not an alias of --store anymore. Therefore, developers remove the code of handling the option:
    // in function run (argv: string[])
    await new Promise((resolve, reject) => {
    setTimeout(() => {
    - if (opts.storePath && !opts.store) {
    - logger.warn(’the ‘store-path‘ config is deprecated. Use ‘store‘ instead.’)
    - opts.store = opts.storePath
    // ‘pnpm install ""‘ is going to be just ‘pnpm install‘
    const cliArgs = cliConf.argv.remain.slice(1).filter(Boolean)
  3. (3)
    Changing the default behavior for unprovided values (203/1,034 cases). Typically, in commit 72bbda7f of npm/cli [^19] https://github.com/npm/cli/commit/72bbda7f (the diff shown below), the default value of local variable “depthToPrint” is set to zero while it is initially “undefined”.
    const { /* ... */, depth, /* ... */ } = npm.flatOptions;
    const depthToPrint = all ? Infinity : (depth || 0);
    The default values can be also in global configurable objects. For example, Renovate puts configurations that affect the behavior in lib/config/options/index.ts, and the modification to it will cause behavioral change.
  4. (4)
    Changing error handling process (42/1,034 cases). One case is changing change the content or format of the error objects. For example, in commit dd6306e3 of octokit/rest.js [^20] https://github.com/octokit/rest.js/commit/dd6306e3, the error message is parsed as a JSON object and the properties will be put into error object (the existing message property will be overwritten). Since developers usually parse error.message, this change will make the code like JSON.parse(error.message) not work.
    // in a Promise chain
    .catch(error => {
    if (error instanceof HttpError) {
    + try {
    + Object.assign(error, JSON.parse(error.message))
    + } catch (_error) {
    + // ignore, see #684
    throw error
    Another case is changing the criteria of detecting errors, e.g, some status values are no longer regarded as success. In the commit 201f189d in angular/angular [^21] https://github.com/angular/angular/commit/201f189d, if the return code of an XHR (XML HTTP Request) is not 200, an error will occur, while before the code change, if the return codes between 200 and 300 (exclusive) were all regarded as success states.
    // Breaking Change in xhr_backend.ts
    let response = new Response(responseOptions);
    + if (isSuccess(status)) {
    responseObserver.next(response);
    responseObserver.complete();
    + return;
    + responseObserver.error(response);
    // in http_util.ts
    + export const isSuccess =
    + (status: number): boolean => (status >= 200 && status < 300);
We also compare the behavioral BCs those in Maven ecosystem. Zhang et al. have studied the behavioral BC types in Maven projects (Zhang et al., 2022). They found that changing execution logic and changing calculation of the output make up over 60% of the total BCs collected by them. They also studied benign changes that will not cause BCs, such as additional/changed/deleted conditions and branches (over 55% of the total benign changes), additional try/catch statements, and assignment revision. However, in our findings about source code patterns of JavaScript behavioral BCs, it is very common to change conditions in “if” statements to remove support for some option values or to change evaluation criteria of some status, although it seems benign. Hence, in JavaScript, the condition changes that involve option/configuration values deserve attention since they are more probable to be BCs.
4.4. RQ4: Why do Developers Make Breaking Changes in NPM Ecosystem?
In this section, we summarize six major motivations behind making BCs by carefully analyzing our collected BCs. We show the connection between BC types and motivations in Figure 3.
Figure 3. Distribution of Motivations in Different Types of Breaking Changes
Reason 1: To reduce code redundancy.
(469 cases) Some source code is no longer needed, hence developers remove them. There are two cases of this reason:
  • To unify the approach of an functionality (21 cases). The functionality can be achieved by other methods, or by using other packages. Therefore, the current implementation can be removed. For example, in commit 1354171 of thi-ng/umbrella, developers removed function “writeFile”, and recommend using “writeFile” provided by another package @thi-ng/rstream-log-file since the third-party function can also achieve the same functionality.
  • To deprecate and remove an outdated functionality (413 cases). Some outdated input (e.g., authentication mechanism) are deprecated thus developers remove their support. For example, in commit ad4f1493 of Rest.js, the support for netrc authentication method was removed while others were retained.
  • To better organize the project (35 cases). Moving part of the functionality out and create a new package will make the project better organized. For example, in commit 0fa2830222222https://github.com/jsdoc/jsdoc/commit/0fa2830 of jsdoc/jsdoc, the package @jsdoc/core was reorganized and some methods in @jsdoc/core moved to a new package @jsdoc/cli.
Table 4. Typical Rename Operations
Category Explanation Number
Add token add non-trivial tokens, e.g., from description to ariaDescription 23
Change token order change orders in the identifier, e.g., secretJsonValue to secretValueFromJson 2
Remove token remove one token from the identifier, e.g., renderBoundElementIndex to boundElementIndex 15
Replace completely replace the token with another token, e.g., signed to trusted 20
Replace token replace one token in identifier, e.g., fromDockerHub to fromDockerRegistry 39
Trivial add is or get prefix, e.g., from failed to isFailed 12
Reason 2: To improve identifier names.
(111 cases) Good identifier names can indicate the functionalities. However, improper identifier names can be a main contributor to technical debt in JavaScript projects. Improper identifier names may come from lack of familiarity with external knowledge and a failure to consider previous code. This reason can be divided into the following categories:
  • To make the identifier name reflect the actual functionality (68 cases). The name of a class, interface or method, etc., should reflect the actual functionality. Therefore, after a code change, the identifier names tend to be modified. For example, in issue 3403 of angular/angular [^23] https://github.com/angular/angular/issue/3403, the developer complaint that “When I first saw the requestCheck() method on ChangeDetectorRef, I assumed it was how I manually run change detection. Others have assumed this as well”. Hence after the issue was reported, the identifier requestCheck is changed to markForCheck.
  • To make the identifier more idiomatic (24 cases). For example, adding prefix is or get to the original name (e.g., from empty to isEmpty), and changing the order of tokens in an identifier (e.g., from listSecretsForRepo to listRepoSecrets).
  • To unify identifier names (12 cases). Similar classes or methods (e.g., components with the same style in a package) should share a similar name. For example, in commit 9bbd2469 of pnpm/pnpm [^24] https://github.com/pnpm/pnpm/commit/9bbd2469, the field localPackages was renamed to workspacePackages. In the dependency @pnpm/resolver-base [^25] This package is part of pnpm, and the Rename BC is located in the package @pnpm/find-workspace-packages, also a part of pnpm., LocalPackages had been renamed to WorkspacePackages, hence to make it consistent, the field was also renamed.
  • To avoid name conflict (5 cases). For example, in commit da19583 of adonisjs/adonis-framework [^26] https://github.com/adonisjs/adonis-framework/commit/da19583, developers renamed Context to HttpContext since Context is a commonly used keyword.
  • To fix typos (2 cases). Simply fixing a typo. For example, renaming “MetricAarmProps” to “MetricAlarmProps”.
We also compared the rename operations in JavaScript (and TypeScript) with those in Java since most Java and JavaScript projects adopt the camelCase naming convention. For JavaScript and TypeScript projects, we show the rename operations in Table 4. For Java projects in Maven ecosystem, we inspect Huang et al.’s dataset on Java API migration (Huang et al., 2021). Interestingly, we find that most method rename BCs are trivial, i.e., only adding or removing prefixes get, is (e.g., from failed to isFailed), and only few classes are renamed with one token replaced added, removed or replaced. Other popular languages like Python, Go and Rust, do not use camelCase convention.
Reason 3: To improve API design.
(939 cases) The quality of API can impact the productivity of programmers, the adoption of APIs, and the quality of dependent code (Myers and Stylos, 2016). This reason can be divided into the following sub-categories:
  • To make API behavior more reasonable (704 cases). The major scenarios of API’s unreasonable behavior include: ① The API design does not conform to a certain standard and users may misuse the API. For example, in the commit 201f189d of angular/angular, after the code change, status codes less than 200 and greater than 299 will cause error, while previously, errors only occurred when network errors. ② The API behavior can easily cause unintended results. For example, in commit 7d4c399 of gajus/eslint-plugin-jsdoc, the behavior is changed unless a new option is set to true. This is to decrease the false positives when capitalized letters on newlines merely represent proper nouns.
  • To make APIs more convenient to use (235 cases). For example, changing the parameter order of a method can be more in line with users’ habits, and using async rather than Promise or callback parameters also provides convenience for downstream developers.
5. Implications
On the basis of the results above, we can yield the following implications for future works:
Automatic naming and renaming techniques can be applied in NPM projects. In Section 4.3, we conclude that the main reason for renaming identifiers is to make identifier name reflect the actual functionality. Therefore, on the one hand, poor identifier name is a noticeable technical debt in JavaScript projects, hence automatic naming can be utilized to provide good identifier names, and on the other hand, in some cases, renaming identifiers is reasonable due to the necessary adaptation to frequent code changes in NPM ecosystem, thus we suggest applying automatic renaming techniques to make identifier names adapt to code evolution.
Several rule-based, static analysis-based, and machine learning-based approaches are proposed to help automatic naming and renaming. For example, the tool proposed by Caprile et al. (Caprile and Tonella, 2000) is a typical rule-based approach that makes unified tokens in an identifier comply to syntax rules, the tool proposed by Feldthaus et al. (Feldthaus and Møller, 2013) is a static analysis-based approach to semi-automatically rename object properties with a focus on related property identifiers [^27] Property identifiers refer to property names of an object, e.g., if option is a parameter, in options.followSymLink, followSymLink is a property identifier., and RefBERT (Liu et al., 2023) is a typical deep learning-based approach that utilizes fune-tuning pre-training model (i.e., RoBERTa) and infers local variable names with contextual token sequences. However, these approaches suffer from two major limitations, and the future works can try to mitigate them:
  • They are not proposed to deal with identifiers that can be directly accessed by downstream developers (e.g., class and interface names), while Rename actions on these identifiers make up the most in JavaScript projects. For instance, the most recent renaming approach RefBERT works in the same way as code completion, i.e., the intra-function context is input to infer local variable names, however, inferring identifier names needs different context knowledge, such as class structure, client invocation code, etc.
  • They are not code-change-aware. Since we have found that many a proportion of Rename BCs are performed to ensure code consistency (Figure 3 in Section 4.3), a renaming technique had better leverage knowledge in prior related code changes to recommend proper names.
Tools for detecting code with similar functionalities are necessary. According to Figure 3, a number of Remove BCs are motivated by approach for unification, i.e., provide a unified API for a specific functionality. However, unlike the casualty changes (e.g., removing a class then the code referring to it should also be removed), the relationship between these co-occurring changes is indirect and not simple to detect. Besides, in our BC extraction process, we find developers make other small changes for better performance, compatibility, and robustness though they are unrelated to BCs. This also to some degree interferes with BC identification. The findings in our investigation can provide the following directions for future research:
  • Investigate how to indicate potential BC actions when one BC has been already performed and recommend similar BC actions to improve code consistency and reduce code redundancy.
  • Identify functionally related code snippets that can be removed together to facilitate clean code.
  • Distinguish non-breaking changes from breaking changes and casualty changes to provide more convenience for developers.
Future research can focus on detecting more types of behavioral BCs. Syntactic BC detection (e.g., Rename, Remove and Change Signature) is somehow trivial and many refactoring detection-based tools can achieve it. For example, APIDiff (Brito et al., 2018), Clirr (Cli, [n. d.]), Japicmp (jap, [n. d.]) and RevAPI (Rev, [n. d.]) can be used for Java projects, PyCompat (Zhang et al., 2020) and Pidiff (pid, [n. d.]) can be applied in for Python libraries, and a variant of APIDiff can be used for JavaScript to characterize refactoring history (Brito et al., 2021). Hence the future research works should focus on behavioral BCs that cannot be simply identified. Existing such approaches for JavaScript are NoRegrets (Mezzetti et al., 2018) and its enhanced version NoRegrets+ (Møller and Torp, 2019). In Section 1, we have highlighted their reliance on test cases in downstream projects. In addition, the two approaches can only detect type-related BCs. Specifically, in the first phase, they run client test cases and monitor the flows of values by program instrumentation to build API models (intuitively, type restriction of APIs), and in the second phase, they rerun the client test cases with updated upstream code to check whether type restrictions are changed during execution. Therefore, they cannot detect non-type-related BCs, e.g., a BC that changes the semantics of the returned string, while the return type is still string. According to Section 4.3, there are five types of behavioral BC actions, where only changing the specifications of return values can directly cause type changes, while other behavioral change may not change object types during execution. Therefore, we recommend future works try to deal with various types of behavioral BCs uncovered by this study.
6. Threats to Validity
Threats to internal validity are mainly related to the clarity of developers’ documentation of breaking changes. Few developers might have not explicitly documented the BCs in commit messages, and commit messages (as well as text in issues and pull requests) might not fully reflect developers’ reason for making BCs. Also, the discussion and documentation related to BCs may not fully reflect developers’ opinions. Second, in our manual analysis, we might still not understand developers’ intentions since many BC commit messages do not clearly explain what and why they make BCs. Additionally, we might regard some changes in code diff as BC actions since they are confusing. To address this potential threat, we have implemented a double-checking process to verify the accuracy of our analysis results.
Threats to construct validity are concerned with the errors during breaking change selection. While checking the compliance to Conventional Commits standard, we simply use the regular expression, which might incorrectly classify some commit messages. Also, in the process of selecting commits for further study (Section 3.3), we might abandon some commits, although they satisfy our selection criteria.
Threats to external validity refer to the generalizability of our study. In our study, while we involve 1,519 BCs extracted from 381 randomly sampled projects from a vast number of popular NPM projects, and these projects cover diverse application fields of JavaScript language, such as utilities, frontend, Web APIs, database, etc., the extracted BCs only account for a small proportion of BCs in the whole NPM ecosystem. Besides, in this study we do not consider less popular projects since they have short development history and it is less probable to yield some patterns from these projects. However, the BC practices in less popular projects may still differ from those in popular projects since they are not so well maintained.
Threats to conclusion validity relate to insufficient high-quality BC commits to support our findings. To mitigate the threats, we try our best to extract breaking changes from commit messages, issues, pull requests and changelogs to construct our BC dataset. We also carefully inspect our collected BCs and retain those related to JavaScript source code changes and explicitly documented by developers, which can help increase the quality of our dataset.
7. Related Work
7.1. Research on NPM Ecosystem
A number of research works have investigated various aspects of the NPM ecosystem, which can help us understand the challenges in NPM ecosystem and provide future research directions. Decan et al. (Decan et al., 2018) studied the impact of vulnerabilities in the NPM ecosystem by analyzing how and when security vulnerabilities are reported and fixed, and to which extent they affect other packages in the NPM ecosystem in the presence of dependency constraints. Liu et al. (Liu et al., 2022) employed a knowledge graph method to characterize vulnerability propagation and evolution in the NPM ecosystem. Cogo et al. (Cogo et al., 2019) conducted a systematic investigation into NPM dependency downgrades. Mujahid et al. (Mujahid et al., 2023) summarized the characteristics of highly-selected projects in NPM. Maeprasart et al. (Maeprasart et al., 2023) illustrated the role of external pull requests in NPM project development. Abdalkareem et al. (Abdalkareem et al., 2017) examined the usage of trivial packages, while Chen et al. (Chen et al., 2021) further uncovered the motivations driving JavaScript developers to publish such packages despite their potential drawbacks. Venturini et al. (Venturini et al., 2023) studied the impact of manifesting breaking changes in NPM, which is the most closely related work to ours. In contrast to their study, our research delves into multiple dimensions of breaking changes, including the affected program elements, the impact on client applications, and the underlying reasons behind these changes. Additionally, we consider the specific language features of JavaScript and TypeScript at a fine-grained level.
7.2. JavaScript Static Analysis
Static analysis is a straightforward way to determine many properties of JavaScript programs and can be used to detect potential bugs and breaking API changes. Some open source tools like ESLint (ESL, [n. d.]) and JSHint (JSH, [n. d.]) support rule-based static analysis for JavaScript projects. They can be used for better code quality, such as making code follow JavaScript programming idioms. Some works proposed static analysis approaches from many aspects. For example, Madsen et al. built event-based call graphs (Madsen et al., 2015) to detect event-related bugs by enhancing the static analysis framework JASI (Kashyap et al., 2014) and TAJS (Jensen et al., 2011). Other works (Park et al., 2015; Sung et al., 2016; Jensen et al., 2011) studied static analysis for JavaScript in the DOM environment. Furthermore, in addressing the limitations of static analysis approaches for JavaScript, Chakraborty et al. presented a technique to supplement missing edges in the JavaScript call graph (Chakraborty et al., 2022). Through their research, they discovered that dynamic property access is a primary factor contributing to low recall in previous static analysis frameworks, typically missing some function invocations. By applying their proposed technique, they were able to improve the recall rate. However, due to the complexity of language features, a static analysis approach cannot cover all language use cases. Also, these static analysis approaches focus on the detection and elimination of ill-formed code, and optimization of the testing process, rather than breaking changes among commits. By applying their proposed technique, they were able to improve the recall rate.
7.3. Breaking Change Detection and Analysis on Other Platforms
Besides JavaScript, a lot of studies focus on detecting and analyzing BCs for other programming languages, especially Java and Python, and their techniques may be learned by BC detection for JavaScript. Brito et al. presented APIDiff (Brito et al., 2018) that can detect syntax-related breaking changes in Maven projects, such as method removal and visibility loss by reusing the refactoring detection tool RefDiff (Silva and Valente, 2017). Some open-source tools can also check Java syntactic breaking changes, such as Clirr (Cli, [n. d.]) and RevAPI (Rev, [n. d.]). For Python language, Du et al. proposed AexPy (Du and Ma, 2022) that can detect similar types of breaking changes like module removal and addition of required parameters, which extends the existing tool PyCompat (Zhang et al., 2020) and Pidiff (pid, [n. d.]). Regarding non-syntactic breaking changes, to the best of our knowledge, Zhang et al. proposed Sembid (Zhang et al., 2022) to detect behavioral BCs by measuring the semantic difference of call graphs between old and new programs: if a code change’s semantic difference is larger than a threshold and not identified as benign change, it is classified as a BC. However, their used semantic diffs reflect the structural changes of programs, while many semantic BCs do not need many of those changes, e.g., only changing one condition, and adding a small change to the output string. Additionally, as the authors have pointed out, Sembid cannot detect high-level BCs such as re-implementation of a method. Therefore, directly adapting Sembid to JavaScript projects is not suitable.
Additionally, with the help of BC detection techniques, many works analyzed the impacts of breaking changes to downstream projects. For example, Jayasuriya et al. (Jayasuriya et al., 2023, 2024) investigated the impacts of breaking changes (especially behavioral breaking changes) to client applications in Maven ecosystem. They concluded with many findings, e.g., 11.58% of the dependency updates contain breaking changes that impact the clients and 2.30% version updates have behavioral breaking changes that impacted client tests. Similar to the findings in Venturini et al.’s study (Venturini et al., 2023), while the breaking changes are only detected in a small number of code changes or version updates, they still cause negative impacts to downstream projects.
8. Conclusion and Future Work
In this study, for better comprehension of breaking changes in the NPM ecosystem and enhancing BC-related tools, we conduct an empirical study to bridge the knowledge gap from three aspects, with our carefully constructed breaking change dataset (1,519 BCs in total) sampled from a large set of popular NPM projects. We find that 97% of BCs detected by regression testing can be covered by developers’ documentation, which proves that extracting BCs from documentation is reasonable. We then summarize the JavaScript and TypeScript specific syntactic BCs and how developers make behavioral BCs, yielding many findings. Besides, we also investigate the reasons behind BCs in JavaScript and conclude with a taxonomy, which extend the previous works on motivations behind breaking changes. Based on our empirical findings, we provide actionable implications for future research, e.g., applying automatic renaming and naming techniques in JavaScript projects, and detecting code with similar functionalities, etc.
In the future, we want to collect additional sources such as online discussions, and employ alternative approaches like dynamic analysis to gain a deeper understanding of breaking changes in the NPM ecosystem. We plan to investigate how to automatically identify the breaking changes in NPM projects. We consider combining static analysis and dynamic analysis techniques to enhance the existing BC detection approaches. We also consider improving BC detection with the help of Large Language Models since they show great performance improvements in many software engineering related tasks and can utilize the semantics in source code.

### v2

@@
 Abstract.
-Code evolution is prevalent in software ecosystems, which can provide many benefits, such as new features, bug fixes, security patches, etc., while still introducing breaking changes (BCs) that make downstream projects fail to work.
+Code evolution is prevalent in software ecosystems, which can provide many benefits, such as new features, bug fixes, security patches, etc., while still introducing breaking changes that make downstream projects fail to work.
 Breaking changes cause a lot of effort to both downstream and upstream developers: downstream developers need to adapt to breaking changes and upstream developers are responsible for identifying and documenting them.
@@
 We construct a dataset of explicitly documented breaking changes from 381 popular NPM projects.
-We find that 93.6% of the detected breaking changes can be covered by developers’ documentation, and about 19% of the breaking changes cannot be detected by regression testing.
+We find that 95.4% of the detected breaking changes can be covered by developers’ documentation, and 19% of the breaking changes cannot be detected by regression testing.
 Then in the process of investigating source code of our collected breaking changes, we yield a taxonomy of JavaScript and TypeScript-specific syntactic breaking changes and a taxonomy of major types of behavioral breaking changes.
-Additionally, we investigate the reasons why developers make breaking changes in NPM and find three major reasons, i.e., to reduce code redundancy, to improve identifier name, and to improve API design, and each category contains several sub-items.
+Additionally, we investigate the reasons why developers make breaking changes in NPM and find three major reasons, i.e., to reduce code redundancy, to improve identifier names, and to improve API design, and each category contains several sub-items.
 We provide actionable implications for future research, e.g., automatic naming and renaming techniques should be applied in JavaScript projects to improve identifier names, future research can try to detect more types of behavioral breaking changes.
-By presenting the implications, we also discuss the weakness of automatic renaming and BC detection approaches, such as the lack of support for public identifiers and various types of breaking changes.
+By presenting the implications, we also discuss the weakness of automatic renaming and breaking change detection approaches, such as the lack of support for public identifiers and various types of breaking changes.
 Breaking Change, NPM, JavaScript, Code Evolution
@@
 For example, renaming frequently used methods or classes can make client projects that depend on the code fail to compile.
-Therefore, downstream developers will be required to make efforts to adapt to such BCs.
-Considering the risks and the effort brought by BCs to downstream developers, many works have been utilized to help developers analyze and measure the impacts of BCs (Mezzetti et al., 2018; Møller and Torp, 2019; Don, [n.
+Therefore, downstream developers will be required to make efforts to adapt to such breaking changes.
+Considering the risks and the effort brought by breaking changes to downstream developers, many works have been utilized to help developers analyze and measure the impacts of breaking changes (Mezzetti et al., 2018; Møller and Torp, 2019; Don, [n.
 d.]).
-However, they relied on test suites in downstream projects, and either detected JavaScript BCs by directly running test suites (Don, [n.
+However, they relied on test suites in downstream projects, and either detected JavaScript breaking changes by directly running test suites (Don, [n.
 d.]), or via API models [^1] Intuitively, an API model defines the type restrictions of an API (Mezzetti et al., 2018).
 generated from dynamic execution of test cases (Mezzetti et al., 2018; Møller and Torp, 2019).
-This leads to the result that BCs observed in prior studies are located in well-known APIs that are popularly used in downstream projects, such as “Client.socket” in socket.io [^2] https://github.com/socketio/socket.io/commit/b73d9be, “Request.prototype.onResponse” in request [^3] https://github.com/request/request/commit/d05b6ba, and each in async.
-However, BCs are reported to upstream projects because they already manifested themselves in downstream test cases or already caused bugs.
-For example, in version 3.0.0 of a famous JavaScript project Lodash, the behavior of function “mixin” was changed and not explicitly marked by upstream developers, and a downstream developer reported this issue by self-constructed test code.
-[^4] https://github.com/lodash/lodash/issues/880 Hence, completely relying on test case execution is not sufficient for understanding BCs.
+This leads to the result that breaking changes observed in prior studies are located in well-known APIs that are popularly used in downstream projects, such as Client.socket in socket.io [^2] https://github.com/socketio/socket.io/commit/b73d9be, Request.prototype.onResponse in request [^3] https://github.com/request/request/commit/d05b6ba, and each in async.
+However, breaking changes are reported to upstream projects because they already manifested themselves in downstream test cases or already caused bugs.
+For example, in version 3.0.0 of a famous JavaScript library lodash, the behavior of function mixin was changed and not explicitly marked by upstream developers, and a downstream developer reported this issue by self-constructed test code.
+[^4] https://github.com/lodash/lodash/issues/880 Hence, completely relying on test case execution is not sufficient for understanding breaking changes.
 We observe that in the NPM ecosystem, a number of projects write commit messages complying with Conventional Commits (Con, [n.
 d.]).
-According to Conventional Commits, “BREAKING CHANGE” tokens are documented in commit messages to indicate that the corresponding commits contain breaking changes.
+According to Conventional Commits, BREAKING CHANGE tokens are documented in commit messages to indicate that the corresponding commits contain breaking changes.
 This is because the core developers of upstream libraries need to ensure the stability of the API and protect the reputation of the upstream projects.
-When a new code change is submitted for review, they carefully inspect it for the presence of BCs, as well as other bugs or defects.
-By leveraging the documented BCs provided by developers, there is a potential opportunity to notify downstream users about these BCs.
-However, developers experience increased pressure within the NPM ecosystem, primarily driven by frequent code changes and a high tolerance for BCs (Bogart et al., 2021, 2016).
-Another possible factor is that different developers may have different perceptions of BCs and some of them may ignore the BCs, leading to unaware BCs and causing unforeseen bugs in downstream projects.
-If we could characterize the large number of documented BCs at source code level, it would assist upstream developers in gaining a deeper understanding of the types of commits that could potentially introduce BCs.
-Furthermore, such categorization would provide valuable insights for researchers to develop tools aimed at detecting BCs in the future.
-To bridge this gap, we conduct an empirical study to better comprehend documented BCs.
-We first select 381 popular JavaScript projects (randomly selected from 35,786 projects with more than 50 GitHub stars in the NPM ecosystem) and clone their associated repositories.
-We then identify BC-related commits admitted by developers by searching for BREAKING CHANGE tokens in commit messages.
-For the 5,242 identified BC commits, we manually select 1,519 commits that 1) containing not too long commit messages, 2) related to JavaScript production code and 3) associated with documentation that can explain the reasons behind the commit.
-These selected BC commits are distributed across 131 distinct projects and the projects can be grouped into six categories according to functionalities and usages, i.e., utilities, frontend projects, development-related tools, database-related tools, plugins and Web API related tools (detailed in Section 3), which makes the BC commits in our projects representative.
-To extract BCs from these BC commits, we build BC type categories by learning from a previous API detection tool for Java (Brito et al., 2018) and adding new JavaScript and TypeScript-specific features.
+When a new code change is submitted for review, they carefully inspect it for the presence of breaking changes, as well as other bugs or defects.
+By leveraging the documented breaking changes provided by developers, there is a potential opportunity to notify downstream users about these breaking changes.
+However, developers experience increased pressure within the NPM ecosystem, primarily driven by frequent code changes and a high tolerance for breaking changes (Bogart et al., 2021, 2016).
+Another possible factor is that different developers may have different perceptions of breaking changes and some of them may ignore the breaking changes, leading to unaware breaking changes that cause unforeseen bugs in downstream projects.
+If we could characterize the large number of documented breaking changes at source code level, it would assist upstream developers in gaining a deeper understanding of the types of commits that could potentially introduce breaking changes.
+Furthermore, such categorization would provide valuable insights for researchers to develop tools aimed at detecting breaking changes in the future.
+To bridge this gap, we conduct an empirical study to better comprehend documented breaking changes.
+We first select popular projects with over 50 GitHub stars and associated links to an existing GitHub repository, resulting in a total of 35,786 repositories.
+We then randomly select 381 repositories (95% confidence level and 5% margin of error) for further investigation.
+We then clone their associated repositories.
+After this, we identify BC-related commits admitted by developers by searching for BREAKING CHANGE tokens in commit messages.
+For the 5,242 identified BC commits, we manually select 1,519 commits that 1) do not contain too long commit messages, 2) are related to JavaScript production code and 3) are associated with documentation that can explain the reasons behind the commit.
+These selected BC commits are distributed across 131 distinct projects and the projects can be grouped into six categories according to functionalities and usages, i.e., utilities, frontend projects, development-related tools, database-related tools, plugins and Web API related tools (detailed in Section 3), which makes the BC-related commits in our projects representative.
+To extract breaking changes from these BC commits, we build breaking change type taxonomy by learning from a previous API detection tool for Java (Brito et al., 2018) and adding new JavaScript and TypeScript-specific features.
 Our taxonomy includes Remove, Rename, Change Signature, Change Behavior types, etc., which is detailed in Section 2.2.
-Through several thematic analyses, we obtain JavaScript and TypeScript specific code features (RQ2 and RQ3) and a taxonomy of reasons behind BCs.
+Through several thematic analyses, we obtain JavaScript and TypeScript specific code features (RQ2 and RQ3) and a taxonomy of reasons behind breaking changes.
 We answer the following research questions in this study:
 RQ1: To what extent do detected breaking changes and documented breaking changes overlap?
-In our work, we collect BCs from developers’ intention, i.e., the explicitly documented BCs.
-We first check whether documented BCs actually break test code.
-Therefore, we do regression testing on our collected BC commits, similar to previous works (Mezzetti et al., 2018; Møller and Torp, 2019; Mostafa et al., 2017; Zhang et al., 2022).
-We find 97% of detected BCs are documented and 81% of the documented BCs can be detected by regression testing.
-The result shows that most detected BCs are well documented, while a proportion of documented BCs cannot be detected by regression testing.
-Since the documented BCs cover most of the detected BCs, hence extracting BCs from documentation (especially commit messages, issues, and pull requests) is reasonable.
+In our work, we collect breaking changes from developers’ intention, i.e., the explicitly documented breaking changes.
+We first check whether documented breaking changes actually break test code.
+Therefore, we do regression testing on our collected BC-related commits, similar to previous works (Mezzetti et al., 2018; Møller and Torp, 2019; Mostafa et al., 2017; Zhang et al., 2022).
+We find 95.4% of detected breaking changes are documented and 81% of the documented breaking changes can be detected by regression testing.
+The result shows that most detected breaking changes are well documented, while a proportion of documented breaking changes cannot be detected by regression testing.
+Since the documented breaking changes cover most of the detected breaking changes, it is reasonable to extract breaking changes from documentation (especially commit messages, issues, and pull requests).
 RQ2: What syntactic breaking changes in the NPM ecosystem are specific to JavaScript and TypeScript?
-We investigate whether there are some BCs related to JavaScript-specific language features.
-Despite the triviality of most syntactic BCs (such as moving classes to another package, removing a field of a class), we find some JavaScript-specific BCs are notable, including two removing operations, i.e., 1) removing default export, 2) removing export of an element in one place, and five types of JavaScript and TypeScript specific signature changes, e.g., parameter changes in configuration objects, switch between callback and Promise, etc.
+We investigate whether there are some breaking changes related to JavaScript-specific language features.
+Despite the triviality of most syntactic breaking changes (such as moving classes to another package, removing a field of a class), we find some JavaScript-specific BCs are notable, including two removing operations, i.e., 1) removing default export, 2) removing export of an element in one place, and five types of JavaScript and TypeScript specific signature changes, e.g., parameter changes in configuration objects, switch between callback and Promise, etc.
 RQ3: How do developers make behavioral breaking changes?
-In this RQ, we investigate how developers perform BCs at source code level.
-Since syntactic BCs can be detected through syntax analysis and code refactoring detection, we mainly focus on behavioral BCs (a.k.a, semantic BCs in some works) that change the internal program logic.
-We find four major types of Change Behavior BC, i.e., 1) changing the specification return values, 2) changing process for some option values, 3) changing default or initial value of variables and 4) changing error handling method.
+In this RQ, we investigate how developers perform breaking changes at source code level.
+Since syntactic breaking changes can be detected through syntax analysis and code refactoring detection, we mainly focus on behavioral breaking changes (a.k.a, semantic breaking changes in some works) that change the internal program logic.
+We find four major types of Change Behavior breaking changes, i.e., 1) changing the specification return values, 2) changing process for some option values, 3) changing default or initial value of variables and 4) changing error handling method.
 RQ4: Why do developers make breaking changes in NPM ecosystem?
-In this RQ, we revisit the rationale of BCs in the NPM ecosystem and find three main factors (each with some sub-factors) that motivate developers to make BCs, i.e., 1) to reduce code redundancy, 2) to improve identifier name, 3) to improve API design, and divide each reason into some sub-items, which extended the previous works (Bogart et al., 2016, 2021; Brito et al., 2020).
-Based on our empirical findings, we provide actionable implications for future research, including 1) automatic renaming deserves much concern in JavaScript projects since poor identifier name is a main contributor to technical debt, 2) automatic tools for ensuring code consistency are necessary, 3) future research can try to detect more types of behavioral BCs.
-By presenting these implications, we also discuss the weakness of automatic renaming and BC detection approaches, such as lacking support for public identifiers and various types of BCs.
+In this RQ, we revisit the rationale of breaking changes in the NPM ecosystem and find three main factors (each with some sub-factors) that motivate developers to make breaking changes, i.e., 1) to reduce code redundancy, 2) to improve identifier names, 3) to improve API design, and divide each reason into some sub-items, which extends the previous works (Bogart et al., 2016, 2021; Brito et al., 2020).
+Based on our empirical findings, we provide actionable implications for future research, including 1) automatic renaming deserves much concern in JavaScript projects since poor identifier naming is a main contributor to technical debt, 2) automatic tools for ensuring code consistency are necessary, 3) future research should strive to detect more types of behavioral breaking changes.
+By presenting these implications, we also discuss the weakness of automatic renaming and breaking change detection approaches, such as lacking support for public identifiers and various types of breaking changes.
 The contributions of our paper are two-fold:
@@
   2. (2)
-    We empirically identify how developers perform BCs at source code level, yielding notable findings like JavaScript and TypeScript specific syntactic BCs and typical actions of behavioral BCs.
-We also investigate why developers perform BCs in the NPM ecosystem.
+    We empirically identify how developers perform breaking changes at source code level, yielding notable findings like JavaScript and TypeScript specific syntactic breaking changes and typical actions of behavioral breaking changes.
+We also investigate why developers perform breaking changes in the NPM ecosystem.
 The remainder of this paper is organized as follows.
@@
 Section 7 reviews the related work and Section 8 concludes our work.
-Data availability. We provide the datasets of our research at https://zenodo.org/records/13219136.
+Data availability. We provide the replication package of our research at https://doi.org/10.5281/zenodo.13927690.
 2. Background
@@
 2.1. Preliminary of Breaking Changes
-Prior to our work, a number of studies have uncovered BCs from many aspects, such as the occurrence of BCs, the impacts of BCs on downstream projects, and motivations for making BCs.
-The research works (Decan and Mens, 2019; Venturini et al., 2023; Raemaekers et al., 2017) have demonstrated the widespread occurrence of BCs in the NPM ecosystem and their effects on client applications.
+Prior to our work, a number of studies have uncovered breaking changes from many aspects, such as the occurrence of breaking changes, the impacts of breaking changes on downstream projects, and motivations for making breaking changes.
+The research works (Decan and Mens, 2019; Venturini et al., 2023; Raemaekers et al., 2017) have demonstrated the widespread occurrence of breaking changes in the NPM ecosystem and their effects on client applications.
 Venturini et al.
@@
 Bogart et al.
-found that coarse-grained motivations for making BCs include requirements and context changes, bugs and new features, rippling effects from upstream changes, and technical debt from postponed changes (Bogart et al., 2016, 2021).
+found that coarse-grained motivations for making breaking changes include requirements and context changes, bugs and new features, rippling effects from upstream changes, and technical debt from postponed changes (Bogart et al., 2016, 2021).
 However, despite much knowledge of BC, to the best of our knowledge, none of the studies precisely defined BC.
-Researchers collected BCs in various forms in previous studies.
-For example, in the evaluation of several breaking change detection tools (Zhang et al., 2022; Mezzetti et al., 2018), researchers collected BCs by running downstream test cases: if a test case of a downstream project could not pass with a newer provider (after the code change), then the code change was identified as a BC.
+Researchers collected breaking changes in various forms in previous studies.
+For example, in the evaluation of several breaking change detection tools (Zhang et al., 2022; Mezzetti et al., 2018), the authors collected breaking changes by running downstream test cases: if a test case of a downstream project could not pass with a newer provider (after the code change), then the code change was identified as a BC.
 However, on the one hand, not all providers are dependent on many client applications.
@@
 In another breaking change detection tool APIDiff (Brito et al., 2018), Brito et al.
-utilized code refactoring detection tool RefDiff (Silva and Valente, 2017) to identify syntax-related BCs (e.g., changing method signature or renaming class), however, as they pointed out, the detected BCs are just breaking change candidates (BCC), since some syntax changes are applied to internal methods and classes, which are not intended for public use.
-To this end, they also asked developers to check whether the detected BCCs are actually BCs (Brito et al., 2020).
-In our study, we adopt Brito et al.’s criteria (Brito et al., 2020), since we also investigate the BCs from developers’ perspective, i.e., how and why they perform BCs.
-Specifically, a BC should 1) be confirmed by developers, 2) be categorized into pre-defined BC types.
-To achieve the first criterion, we identify BCs by searching explicit BC declarations in the documentation (usually commit messages).
-To achieve the second criterion, we define several BC types for JavaScript and TypeScript on the basis of Brito et al.’s Java BC types (Brito et al., 2018) and JavaScript development experience.
-We detail the BC types in Section 2.2 and describe how to extract BCs in Section 3.3.
+utilized the code refactoring detection tool RefDiff (Silva and Valente, 2017) to identify syntax-related breaking changes (e.g., changing method signature or renaming class).
+However, as they pointed out, the detected breaking changes are just breaking change candidates (BCC), since some syntax changes are applied to internal methods and classes, which are not intended for public use.
+To this end, they also asked developers to check whether the detected BCCs are actually breaking changes (Brito et al., 2020).
+In our study, we adopt Brito et al.’s criteria (Brito et al., 2020), since we also investigate the breaking changes from developers’ perspective, i.e., how and why they perform breaking changes.
+Specifically, a breaking change should 1) be confirmed by developers, 2) be categorized into pre-defined BC types.
+To achieve the first criterion, we identify breaking changes by searching explicit BC declarations in the documentation (usually commit messages).
+To achieve the second criterion, we define several breaking change types for JavaScript and TypeScript on the basis of Brito et al.’s Java BC types (Brito et al., 2018) and JavaScript development experience.
+We detail the BC types in Section 2.2 and describe how to extract breaking changes in Section 3.3.
 2.2. Types of Breaking Changes in NPM Projects
-Table 1. Syntactic Breaking Change Types Adopted from Previous Works
+Table 1. Breaking Change Types Adopted from Previous Works
 Type Supported Elements
@@
 Change Signature method, field
+Change Behavior the content of method, field, constant and variable
 Existing BC detection tools for Java, typically APIDiff (Brito et al., 2018), support many syntactic and object-oriented programming (OOP) related BC types, such as remove classes and push down fields.
 Since OOP features are also supported in JavaScript (since ECMAScript 2015 (ECM, [n.
-d.]), class can be directly used while legacy code must leverage prototype mechanism (inh, [n.
-d.])), we learn from the BC types for Java language and make them adapt to JavaScript and TypeScript (a superset of JavaScript, in practice many source code files are written in TypeScript and will be compiled into JavaScript files) by retaining available BC actions in JavaScript and adding more BCs which are not considered previously.
-We then construct a category of JavaScript and TypeScript BCs (shown in Table 1), where interface, enum and type are only available in TypeScript.
+d.]), class can be directly used while legacy code must leverage the prototype mechanism (inh, [n.
+d.])), we learn from the breaking change types for the Java language and adapt them to JavaScript and TypeScript (a superset of JavaScript, in practice many source code files are written in TypeScript and will be compiled into JavaScript files) by retaining available BC actions in JavaScript and adding more breaking change types which are not considered previously.
+We then construct a taxonomy of JavaScript and TypeScript breaking changes (shown in Section 1), where interface, enum and type are only available in TypeScript.
 We explain each type as follows:
   1. (1)
-    Rename BC refers to identifier changes of public code elements.
-Rename BCs can be applied to almost all code elements, such as exported classes, interfaces, and enums, as well as methods and fields in exported classes and interfaces.
+    Rename refers to identifier changes of public code elements.
+Rename BCs can be applied to many code elements, such as exported classes, interfaces, and enums, as well as methods and fields in exported classes and interfaces.
   2. (2)
-    Remove BC refers to removal of public code elements.
-Remove BC can be applied to almost all code elements.
+    Remove refers to removal of public code elements.
+Remove BC can be applied to many code elements.
 After a Remove BC, the affected public element is not accessible by downstream applications, e.g., an exported class is completely removed, or no longer marked as exported.
   3. (3)
-    Move BC can be applied to all APIs supported by Rename BCs. For example, a class is moved to another package since its functionality is more related to that package.
+    Move can be applied to all APIs supported by Rename BCs. For example, a class is moved to another package since its functionality is more related to that package.
   4. (4)
@@
 By contrast, normal “Remove” actions do not copy the content of the removed method to another existing method.
-However, since inline is a very common refactoring operation in Object-oriented languages, we still regard it as a separate type of BC, like previous works (Brito et al., 2018; Silva and Valente, 2017).
+However, since inline is a very common refactoring operation in object-oriented languages, we still regard it as a separate breaking change type, like previous works (Brito et al., 2018; Silva and Valente, 2017).
   5. (5)
@@
   6. (6)
-    Change Signature BC refers to the modification of method signatures except for directly changing method names.
+    Change Signature refers to the modification of method signatures except for directly changing method names.
 Adding or removing modifiers (such as “static”, “private”) and changing the parameters are typical refactoring actions of this type.
-Besides syntactic BCs, Change Behavior BC (also called semantic BC in some works (Mezzetti et al., 2018; Zhang et al., 2022)) refers to those operations changing the internal behavior (e.g., programmatic logic in methods, and content of global constants that can be used by some methods) rather than syntactic elements, such as changing class names and adding required parameters.
-In other words, after a behavioral BC, the affected classes (or methods, interfaces, etc.) can be invoked as the same way as before.
-In our work, we focus on BCs in JavaScript source code, hence we do not take BCs in non-source code files into consideration, typically CSS, Markdown, HTML, package.json files and configuration files related to TypeScript, Webpack and ESLint, etc.
+  7. (7)
+    Change Behavior (also called semantic breaking changes in some works (Mezzetti et al., 2018; Zhang et al., 2022)) refers to those operations changing the internal behavior (e.g., programmatic logic in methods, and content of global constants that can be used by some methods) rather than syntactic elements, such as changing class names and adding required parameters.
+In other words, after a behavioral breaking change is performed, the affected classes (or methods, interfaces, etc.) can be invoked as the same way as before.
+In our work, we focus on breaking changes in JavaScript source code, hence we do not take breaking changes in non-source code files into consideration, typically CSS, Markdown, HTML, package.json files and configuration files related to TypeScript, Webpack and ESLint, etc.
 3. Methodology
 3.1. Data Collection
-Due to the vast number of JavaScript projects in NPM, our study focuses on selecting the most popular projects to ensure that the collected BCs are representative.
+Due to the vast number of JavaScript projects in NPM, our study focuses on selecting the most popular projects to ensure that the collected breaking changes are representative.
 To do so, we utilize the Libraries.io open source repository and dependency metadata provided by Reid et al.
 (Reid, 2020) to retrieve the most popular JavaScript and TypeScript projects since Libraries.io is commonly used in prior studies (He et al., 2021; Decan et al., 2019, 2018).
-We first select popular projects having more than 50 GitHub stars and associated with an existent repository (deprecated and unmaintained repositories are not considered), which yields 35,786 repositories.
+We first select popular projects having more than 50 GitHub stars and associated with an existent GitHub repository, which yields 35,786 repositories.
 We then randomly select 381 repositories (95% confidence level and 5% margin of error) for further investigation.
-Since we consider the BCs that are explicitly confirmed by developers, we try to extract BCs from developers’ documentation, typically commit messages, changelogs, issues and pull request text on GitHub.
+Since we consider the breaking changes that are explicitly confirmed by developers, we try to extract breaking changes from developers’ documentation, typically commit messages, changelogs, issues and pull request text on GitHub.
 We notice that Conventional Commits (Con, [n.
@@
 BREAKING CHANGE: compiler option has been renamed to runtimeCompiler
-We first check whether Conventional Commits are widely used, analyzing the commits from the 381 projects we used and find that We find that 360 out of 381 repositories contain commits that conform to Conventional Commits and 195 projects have over 80% commits that follow Conventional Commits.
+We first check whether Conventional Commits are widely used, analyzing the commits from the 381 projects we used.
+We find that 360 out of 381 repositories contain commits that conform to Conventional Commits and 198 projects have over 80% commits that follow Conventional Commits.
 We use the regular expression below to check Conventional Commit compliance:
 […]
-This indicates that Conventional Commit is widely used and we can use it as a simple way to extract plenty of BCs.
+Here fix, feat, chore, etc., are the defined or recommended scope tokens in Conventional Commits specification.
+This indicates that Conventional Commits specification is widely used and we can use it as a simple way to extract plenty of breaking changes.
 Considering that many projects do not follow this standard, we also try our best to extract BC commits from changelogs of each project (if developers have declared), issues and pull requests on GitHub.
-For example, the project tj/commander.js does not use Conventional Commits, then we consult the mentioned pull request IDs (and commits related to the PR) in CHANGELOG.md file [^5] https://github.com/tj/commander.js/blob/master/CHANGELOG.md, shown in Figure 1.
+A changelog summarizes the changes from the last version release, and may mention multiple breaking changes.
+By analyzing the changelog, we can link the mentioned breaking changes to corresponding commits.
+For example, the project tj/commander.js does not use Conventional Commits, then we consult the mentioned pull request IDs in CHANGELOG.md file [^5] https://github.com/tj/commander.js/blob/master/CHANGELOG.md (shown in Figure 1, where four breaking changes are mentioned), and we also analyze the content of the pull requests to obtain the commits related to the breaking changes.
 In this way, we obtain 5,242 commits.
-Figure 1. Screenshot of the Changelog of tj/commander.js (Breaking Changes Are Highlighted)
+Figure 1. Screenshot of the Changelog of commander.js 8.0.0 (The four highlighted breaking changes can be linked to related issues and commits)
 Then two of the authors manually inspect them and remove some of them that are not related to JavaScript production code (e.g., only modifying package.json, Markdown documentation, build scripts, test cases) or contain very long commit messages (over 10 lines) that are difficult to understand.
@@
 Additionally, we do not include the commits that contain multiple breaking change declarations.
-Specifically, we remove the 965 commits that only updated dependency versions, 456 commits that only changed minimum versions of supported platforms, and 244 commits that only changed build scripts or JSON configuration files.
-We also remove 865 commits that contain over 10 line commit messages or multiple BREAKING CHANGE declarations.
+Specifically, we remove 1,826 commits that contain over 10 line commit messages or multiple BREAKING CHANGE declarations and 692 commits that are not JavaScript code changes.
 Finally, we retain 2,724 candidate commits.
 3.2. Regression Testing
-We use regression testing to detect BCs in our selected projects (381 in total) to check:
+We use regression testing to detect breaking changes in our selected projects (381 in total) to check:
   1. (1)
-    what percentage of detected BCs are well documented,
+    what percentage of detected breaking changes are well documented,
   2. (2)
-    what percentage of documented BCs can be detected by regression testing.
-To check ①, due to the number of the total commits is too large (1,455,003), we first sample 16,453 commits for further analysis (99% confidence level and 1% margin of error).
+    what percentage of documented breaking changes can be detected by regression testing.
+To check ①, due to the number of the total commits being too large (1,005,344), we first sample 16,371 commits for further analysis (99% confidence level and 1% margin of error).
 We then remove the commits in these projects that only contain non-JavaScript code changes (e.g., test code changes, Markdown and HTML documentation changes, dependency updates in package.json).
@@
     If a detected BC commit is documented with BREAKING CHANGE, or can be found in changelogs, issues, or pull requests, we regard the commit as a documented BC commit.
-We use the test cases written by developers rather than dependents since 1) we have found that all of the 381 projects provide test suites and developers’ test cases can better express the intended usage of APIs (Dinella et al., 2022), while many of the 381 projects do not have plenty of dependents, 2) some projects are not intended for pragmatic use, especially CLI tools in our projects such as npm/cli, therefore no clients access the APIs in these projects.
-To check ②, we apply the steps described above to the 2,724 BC commits from Section 3.1.
+We use the test cases written by developers rather than dependents since 1) we have found that 363 out of the 381 projects provide test suites, and developers’ test cases can better express the intended usage of APIs (Dinella et al., 2022), while many of the 381 projects do not have plenty of dependents, 2) some projects are not intended for pragmatic use, especially CLI tools in our projects such as npm/cli, therefore no clients access the APIs in these projects.
+To check ②, we also apply the steps described above to the 2,724 BC-related commits from Section 3.1.
 3.3. Breaking Change Selection and Analysis
 3.3.1. Breaking Change Selection and Categorization
-We select feasible BCs for further research from the 2,724 candidate BC commits and categorize them into types described in Section 2.2.
+We select feasible breaking changes for further research from the 2,724 candidate BC commits and categorize them into types described in Section 2.2.
 Since we need to understand what a commit changed and why developers made it, we require the commit message and other documentation of a commit (including issues, pull requests and changelogs) to contain some explanation of why developers made that commit.
 We here reuse Tian et al.’s taxonomy of why information in commit messages (Tian et al., 2022).
-Typical why information include:
+Typical why information includes:
   1. (1)
@@
   2. (2)
-    Requirements Illustration: about the usage need, out-of-date statement, etc.
+    Requirements Illustration: about the usage need, out-of-date statements, etc.
   3. (3)
@@
 Figure 2. Categories of Our Investigated Breaking Changes
-Figure 2 presents the percentages of different types of BCs.
-It can be obviously seen that Change Behavior BCs make up the largest proportion of all BC types.
-By contrast, Move BCs make up the smallest percentage.
+Figure 2 presents the percentages of different types of breaking changes.
+It can be obviously seen that Change Behavior BCs make up the largest proportion of all BC types (68.1%).
+By contrast, Move and Inline BCs make up the smallest percentage (1.8% and 0.1% respectively).
 In our further investigation, we mainly focus on Change Behavior BCs.
@@
 Change Behavior 44/45 41/45 25/25 4/4 9/9 3/3 130/131
-The projects that contain BCs in our study serve various types of functionalities, including command-line interface utilities, REST API SDKs and Web frontend frameworks.
+The projects that contain breaking changes in our study serve various types of functionalities, including command-line interface utilities, REST API SDKs and Web frontend frameworks.
 They can be grouped into six categories (shown in Table 2).
-For example, pnpm/pnpm [^7] https://github.com/pnpm/pnpm is a alternative CLI tool for official NPM implementation, renovatebot/renovate provides automatic project build and aws/aws-cdk [^8] https://github.com/aws/aws-cdk wraps the AWS cloud APIs, gajus/eslint-plugin-flowtype [^9] https://github.com/gajus/eslint-plugin-flowtype is a ESLint plugin, and async/async [^10] https://github.com/async/async is a fundamental utility of asynchronous programming.
-Table 3 shows the number of involved projects of each type of BCs.
+For example, pnpm/pnpm [^7] https://github.com/pnpm/pnpm is a alternative CLI tool for official NPM implementation, renovatebot/renovate provides automatic project build and aws/aws-cdk [^8] https://github.com/aws/aws-cdk wraps the AWS cloud APIs, gajus/eslint-plugin-flowtype [^9] https://github.com/gajus/eslint-plugin-flowtype is an ESLint plugin, and async/async [^10] https://github.com/async/async is a fundamental utility of asynchronous programming.
+Table 3 shows the number of involved projects for each type of breaking changes.
 3.3.3. Breaking Change Labeling
-For the 1,519 remaining commits, two of the authors conduct two thematic analyses to analyze ① source code features (for RQ2 and RQ3) and ② reasons behind the commit (for RQ4).
-We use the guideline recommended by Cruzes et al.
-(Cruzes and Dyba, 2011).
+For the 1,519 remaining commits, two of the authors conducted two thematic analyses to analyze ① source code features (for RQ2 and RQ3) and ② reasons behind the commit (for RQ4).
+The guideline recommended by Cruzes et al.
+(Cruzes and Dyba, 2011) were used.
 The guideline for the two analyses is shown below.
-Two of the authors independently perform the steps above, and they hold a series of meetings to solve the agreements.
+Two of the authors independently performed the steps above, and they held a series of meetings to solve the agreements.
   1. (1)
-    We read the related documentation, especially commit messages carefully to understand what the developers want to express.
+    The authors read the related documentation, especially commit messages carefully to understand what the developers want to express.
   2. (2)
-    We read the documentation again and generate phrases as initial codes that describe the source code features, how developers make breaking changes and the reasons behind the commit.
+    The authors read the documentation again and generate phrases as initial codes that describe the source code features, how developers make breaking changes and the reasons behind the commit.
   3. (3)
-    We aggregate the codes with similar meanings and generate a theme name to describe each cluster.
+    The authors aggregated the codes with similar meanings and generate a theme name to describe each cluster.
   4. (4)
-    We then review all themes and try to merge the themes with similar semantics, or make the similar themes become sub-items of a new theme.
-After we obtain all themes, we assign possible themes to each BC.
-Note that for simplicity, we assign each retained BC with one reason, which is consistent with the previous works on motivations behind breaking changes (Brito et al., 2020).
-In this process, we held several meetings to resolve the disagreements on the reason behind each commit since some BC commits might have multiple potential reasons.
+    The authors then reviewed all themes and try to merge the themes with similar semantics, or made the similar themes become sub-items of a new theme.
+After obtaining all themes, the authors assigned possible themes to each breaking change.
+Note that for simplicity, we assigned each retained breaking change with one reason, which is consistent with the previous works on motivations behind breaking changes (Brito et al., 2020).
+In this process, the authors held several meetings to resolve the disagreements on the reason behind each commit since the reasons behind some BC commits are difficult to determine.
 For example, in the commit 2713380 of reactivex/rxjs [^11] https://github.com/reactivex/rxjs/commit/2713380, developers renamed inspect to audit, the reasons behind this rename operation was presented in Issue #1505121212https://github.com/reactivex/rxjs/issues/1505 and #1387131313https://github.com/reactivex/rxjs/issues/1387.
 Developers mentioned two possible reasons, i.e., name collision (since inspect conflicts with util.inspect provided by Node.js [^14] See explanation in Issue 1387 of Rxjs.) and synonym replacement (since audit is synonymous to inspect and it is less confusing).
-By carefully reading and understanding the related issues, two of the authors reach an agreement and assign the reason “to avoid name conflict” to this BC, because if inspect were not used in Node.js native objects, the identifier inspect is still acceptable and does not need to be renamed.
-In our dataset, the authors have very few disagreements due to multiple potential reasons.
+By carefully reading and understanding the related issues, two of the authors reach an agreement and assign the reason “to avoid name conflict” to this breaking change, because if inspect were not used in Node.js native objects, the identifier inspect is still acceptable and does not need to be renamed.
+In the process of assigning the reason to the breaking changes, the authors have very few disagreements (the Cohen’s Kappa value 0.94).
 4. Results
@@
 4.1. RQ1: To What Extent do Detected Breaking Changes and Documented Breaking Changes Overlap?
-From 16,453 commits used in regression testing in Section 3.2, we detect 295 BC commits (1.8%), and 276 of them (93.6%) are documented by developers.
+From 16,371 commits used in regression testing in Section 3.2, we detect 173 BC commits (1.0%), 95.4% of them (165) are documented by developers.
 And for 2,724 BC commits, 2,206 (81%) can be detected by regression testing.
-The results illustrate that most detected BCs are well documented, while a proportion (19%) of documented BCs cannot be detected via regression testing since developers might forget to write test cases or current test cases are too general to cover the code modifications.
-The undocumented BC commits are possibly due to ① the commit is made too long ago when Conventional Commits was not proposed, and ② the project does not comply to Conventional Commits.
-And the main reasons that regression testing cannot detect BCs is that the BC behavior can only be triggered by external conditions (e.g., network error), hence it is difficult to simulate such a situation.
-For example, in commit 61e7a81a of octokit/rest.js, the “followRedirects” option is no longer supported.
+The results illustrate that most detected breaking changes are well documented, while a proportion (19%) of documented breaking changes cannot be detected via regression testing since developers might forget to write test cases or current test cases are too general to cover the code modifications.
+The undocumented breaking change commits are possibly due to ① the commit is made too long ago when Conventional Commits was not proposed, and ② the project does not comply to Conventional Commits.
+And the main reasons that regression testing cannot detect breaking changes is that the BC behavior can only be triggered by external conditions (e.g., network error), hence it is difficult to simulate such a situation.
+For example, in commit 61e7a81a of octokit/rest.js, the followRedirects option is no longer supported.
 This BC can only manifest itself when the HTTP server returns a status code between 301 and 307.
@@
 Developers might forget to write test cases and current test cases might be too general to cover the code modifications, hence automatic test case generation may be useful.
-For the 19% of BCs that cannot be detected, since developers in popular projects have rich development experience, we believe those BCs are trustable and worth attention.
-Although some BCs currently do not affect many downstream projects, they should still be pointed out since in some cases downstream developers may write code that invokes the broken API, then the downstream developers can confirm the breaking changes quickly by looking up commit messages.
-For example, in commit 669592d of Socket.io [^15] https://github.com/socketio/socket.io/commit/669592d contains a breaking change that removed Socket#binary method.
+For the 19% of breaking changes that cannot be detected, since developers in popular projects have rich development experience, we believe those breaking changes are trustable and worth attention.
+Although some breaking changes currently do not affect many downstream projects, they should still be pointed out since in some cases downstream developers may write code that invokes the broken API, then the downstream developers can confirm the breaking changes quickly by looking up commit messages.
+For example, in commit 669592d of socketio/socket.io [^15] https://github.com/socketio/socket.io/commit/669592d contains a breaking change that removed Socket#binary method.
 It is not detected when we run regression testing.
 However, after several months, one developer asked for an alternative method of Socket#binary [^16] https://github.com/socketio/socket.io/discussions/3826.
-This indicates that although an API in a BC is not frequently used, and currently no test cases can trigger it, the API is possibly invoked at a certain time.
+This indicates that although an API in a breaking change is not frequently used, and currently no test cases can trigger it, the API is possibly invoked at a certain time.
 4.2. RQ2: What Syntactic Breaking Changes in NPM Ecosystem Are Specific to JavaScript and TypeScript?
-While most code-level actions in BCs are also available in other programming languages, such as changing parameter orders and removing classes, etc., we highlight the noticeable syntactic BC actions in our investigated projects.
+While most code-level actions in breaking changes are also available in other programming languages, such as changing parameter orders and removing classes, etc., we highlight the noticeable syntactic BC actions in our investigated projects.
 4.2.1. JavaScript-specific remove operations
@@
 like other OOP languages, but also just remove them from exported list and put them into internal source code files.
-We present the JavaScript-specific remove operations in our collected BC commits as follows:
-  1. (1)
-    Remove a default export (6/211 cases).
+We present the JavaScript-specific Remove operations in our collected breaking changes as follows:
+[…]
+Remove a default export (6 cases).
+ECMAScript 6 supports default export functionality [^17] https://developer.mozilla.org/en-US/docs/Web/JavaScript/Reference/Statements/export.
 For example, in commit 102e4b0 of nodkz/mongodb-memory-server, developers simply removed default export of the module util.
 Therefore, the code import generateDbName from 'util' will not work.
-  2. (2)
-    Remove an export position (17/211 cases).
-For example in commit cfbfaac of reactivex/rxjs, the types in rxjs/interfaces modules are no longer accessible and users must use them by importing them from the main module.
+[…]
+Remove an export position (17 cases).
+For example, in commit cfbfaac of reactivex/rxjs, the types in rxjs/interfaces modules are no longer accessible and users must use them by importing them from the main module.
 Therefore, users can import UnaryFunction with the following two methods, while after the code change, they can only import UnaryFunction using the second form.
-    import { UnaryFunction } from ’rxjs/interfaces’;
-    import { UnaryFunction } from ’rxjs’;
+1import { UnaryFunction } from ’rxjs/interfaces’;
+2import { UnaryFunction } from ’rxjs’;
 4.2.2. JavaScript-specific signature changes
@@
 We present the categories as follows:
-  1. (1)
-    About parameters in configuration objects (15/134 cases).
+[…]
+About parameters in configuration objects (15 cases).
 In JavaScript, many properties can be packaged in one configuration object.
-Considering the diff shown below, the second parameter of the method “userinfo” is an object, and this function only extracts four properties in this object using destructing assignment syntax (des, [n.
+Considering the diff shown below, the second parameter of the method userinfo is an object, and this function only extracts four properties in this object using destructuring assignment syntax (des, [n.
 d.]).
-After the code change, the “verb” property is renamed to “method”, since “GET” is an HTTP method.
+After the code change, the verb property is renamed to method, since GET is an HTTP method.
 Correspondingly, the internal code related to this property is also changed.
 async userinfo(accessToken,
-    - verb=’GET’, via=’header’, tokenType, params,
-    + method=’GET’, via=’header’, tokenType, params,
-  2. (2)
-    About this parameter (6/134 cases).
-In the old implementation, the method accepts a callable object (typically a function) and this, and binds the function to this171717bind is a native function in JavaScript.
+- { verb=’GET’, via=’header’, tokenType, params } = {}
++ { method=’GET’, via=’header’, tokenType, params } = {}
+[…]
+About this parameter (6 cases).
+In the old implementation, the method accepts a callable object (typically a function) and this, and binds the function to this181818bind is a native function in JavaScript.
 See https://developer.mozilla.org/en-US/docs/Web/JavaScript/Reference/Global_Objects/Function/bind.
@@
 predicate: (value: T, index: number, source: Observable<T>) => boolean,
-    - thisArg?: any
+- thisArg?: any
 ): OperatorFunction<T, number> {
-    - return operate(createFind(predicate, thisArg, ’index’));
+- return operate(createFind(predicate, thisArg, ’index’));
 + return operate(createFind(predicate, undefined, ’index’));
-  3. (3)
-    About sync and async (20/134 cases).
+[…]
+About sync and async (20 cases).
 In JavaScript, asynchronous functions can be implemented by using callback arguments and Promise objects.
 Legacy JavaScript asynchronous functions are designed with callback-style interface, while new implementations often use async directly.
-For example, in the commit 8c3cecae of Mongoose, developers provided multiple signatures of an API before the code change, e.g., createIndex have the following function signature:
+For example, in the commit 8c3cecae of automattic/mongoose, developers provided multiple signatures of an API before the code change, e.g., createIndex have the following function signature:
 1createIndexes(options: mongodb.CreateIndexesOptions, callback: CallbackWithoutResult): void;
@@
 After the code change in this commit, the signature 1 and 2 were removed. The code diff below shows this pattern:
-    - SomeFunction(param, (returnValue, err) => {
-    - // Some code about returnValue
+- SomeFunction(param, (returnValue, err) => {
+- // Some code about returnValue
 + let returnValue = await SomeFunction(param);
@@
 Another case is converting synchronous to asynchronous (or vice versa), e.g., simply adding async modifier to the function signature.
-  4. (4)
-    About undefined and null (8/134 cases).
+[…]
+About undefined and null (8 cases).
 The keyword null represents an intentional empty value, while undefined occurs in accessing uninitialized variables or non-existent object properties, and should not be explicitly assigned to an object.
 The diff below shows this pattern:
-    - public get(key: Key): Value | null {
-    - return this.collection.get(key) || null;
+- public get(key: Key): Value | null {
+- return this.collection.get(key) || null;
 + public get(key: Key): Value | undefined {
@@
 In this case, developers cannot represent the change in method signature, therefore, we classify the case as behavioral change.
-  5. (5)
-    Switch between “required” and “optional” of fields in method or interface declarations (3/134 cases).
+[…]
+Switch between “required” and “optional” of fields in method or interface declarations (3 cases).
 TypeScript script projects and part of JavaScript projects use d.ts files to declare interface and method signature.
@@
 this: T,
-    - urlParams?: Partial<AbstractInstanceType<T>>,
+- urlParams?: Partial<AbstractInstanceType<T>>,
 + urlParams: Partial<AbstractInstanceType<T>>,
 ): string {
-    - if (urlParams) {
+- if (urlParams) {
+- if (
+- Object.prototype.hasOwnProperty.call(urlParams, ’url’) &&
 4.3. RQ3: How do Developers Make Behavioral Breaking Change?
 We identify four main types of behavioral changes in our collected JavaScript projects of NPM ecosystem. Each type is presented as follows:
-  1. (1)
-    Changing the specification of return values (79/1,034 cases).
+[…]
+Changing the specification of return values (79 cases).
 One typical case is that the type of returned objects are changed.
-For example, in commit b99f6d3 of automattic/mongoose, MongooseArray.map() returned a plain JavaScript array rather than a headless Mongoose array:
+For example, in commit b99f6d3 of automattic/mongoose, the method MongooseArray.map() returned a plain JavaScript array rather than a headless Mongoose array:
 map() {
-    - const ret = super.map.apply(this, arguments);
-    - ret[arraySchemaSymbol] = null;
-    - ret[arrayPathSymbol] = null;
-    - ret[arrayParentSymbol] = null;
+- const ret = super.map.apply(this, arguments);
+- ret[arraySchemaSymbol] = null;
+- ret[arrayPathSymbol] = null;
+- ret[arrayParentSymbol] = null;
 + const copy = [].concat(this);
-    - return ret;
+- return ret;
 + return Array.prototype.map.apply(copy, arguments);
-  2. (2)
-    Changing the process of some options (231/1,034 cases).
+[…]
+Changing the process of some options (231 cases).
 Suppose that one possible value of an option (can be in JSON configuration or parameters) is not supported, then the code that gets properties from the option will be changed, and the corresponding branch of a specific option value will be removed.
@@
 if (!options.type ||
-    - ’basic|oauth|client|token|integration|netrc’.indexOf(options.type) $=$ -1
+- ’basic|oauth|client|token|integration|netrc’.indexOf(options.type) $=$ -1
 + ’basic|oauth|client|token|integration’.indexOf(options.type) $=$ -1
 throw new Error("Invalid authentication type, must be" +
-    - ’basic’, ’integration’, ’oauth’, ’client’ or ’netrc’"
-    + ’basic’, ’integration’, ’oauth’, or ’client’"
+- ’basic’, ’integration’, ’oauth’, ’client’ or ’netrc’")
++ ’basic’, ’integration’, ’oauth’, or ’client’")
 In CLI tool projects, developers often change the process of some options to remove or change the purpose of an option.
 For example, in commit f6fd0c3 of pnpm/pnpm, the option --store-path is not an alias of --store anymore.
-Therefore, developers remove the code of handling the option:
+Therefore, developers remove the code of handling the option, shown as follows:
 // in function run (argv: string[])
@@
 setTimeout(() => {
-    - if (opts.storePath && !opts.store) {
-    - logger.warn(’the ‘store-path‘ config is deprecated. Use ‘store‘ instead.’)
-    - opts.store = opts.storePath
+- if (opts.storePath && !opts.store) {
+- logger.warn(’the store-path config is deprecated.’)
+- opts.store = opts.storePath
 // ‘pnpm install ""‘ is going to be just ‘pnpm install‘
 const cliArgs = cliConf.argv.remain.slice(1).filter(Boolean)
-  3. (3)
-    Changing the default behavior for unprovided values (203/1,034 cases).
-Typically, in commit 72bbda7f of npm/cli [^19] https://github.com/npm/cli/commit/72bbda7f (the diff shown below), the default value of local variable “depthToPrint” is set to zero while it is initially “undefined”.
-    const { /* ... */, depth, /* ... */ } = npm.flatOptions;
-    const depthToPrint = all ? Infinity : (depth || 0);
+// more code ...
+[…]
+Changing the default behavior for unprovided values (203 cases).
+The arguments of a function can often affect the behavior.
+When the value of a parameter is not provided (i.e., being undefined), developers may design special program logic to deal with this case, or make the parametre still remain undefined.
+Sometimes the default program logic of dealing with such case can be changed.
+As an example, in commit 72bbda7f of npm/cli [^20] https://github.com/npm/cli/commit/72bbda7f, the default value of local variable depthToPrint is set to zero while it is initially undefined.
+After this change, the code is as follows:
+1const { /* ... */, depth, /* ... */ } = npm.flatOptions;
+2const depthToPrint = all ? Infinity : (depth || 0);
+3// more code
 The default values can be also in global configurable objects.
 For example, Renovate puts configurations that affect the behavior in lib/config/options/index.ts, and the modification to it will cause behavioral change.
-  4. (4)
-    Changing error handling process (42/1,034 cases).
+[…]
+Changing error handling process (42 cases).
 One case is changing change the content or format of the error objects.
@@
 Since developers usually parse error.message, this change will make the code like JSON.parse(error.message) not work.
+The code diff below shows this change:
 // in a Promise chain
+.then(data => { /* ... */ })
 .catch(error => {
@@
 throw error
-    Another case is changing the criteria of detecting errors, e.g, some status values are no longer regarded as success.
+Another case of this category is changing the criteria of detecting errors, e.g, some status values are no longer regarded as success.
 In the commit 201f189d in angular/angular [^22] https://github.com/angular/angular/commit/201f189d, if the return code of an XHR (XML HTTP Request) is not 200, an error will occur, while before the code change, if the return codes between 200 and 300 (exclusive) were all regarded as success states.
-    // Breaking Change in xhr_backend.ts
+The breaking change in xhr_backend.ts is shown as follows:
 let response = new Response(responseOptions);
@@
 + responseObserver.error(response);
-    // in http_util.ts
+The code in http_util.ts that contains the new anonymous function isSuccess:
 + export const isSuccess =
 + (status: number): boolean => (status >= 200 && status < 300);
-We also compare the behavioral BCs those in Maven ecosystem.
+We also compare the behavioral breaking changes to those in the Maven ecosystem.
 Zhang et al.
-have studied the behavioral BC types in Maven projects (Zhang et al., 2022).
-They found that changing execution logic and changing calculation of the output make up over 60% of the total BCs collected by them.
-They also studied benign changes that will not cause BCs, such as additional/changed/deleted conditions and branches (over 55% of the total benign changes), additional try/catch statements, and assignment revision.
-However, in our findings about source code patterns of JavaScript behavioral BCs, it is very common to change conditions in “if” statements to remove support for some option values or to change evaluation criteria of some status, although it seems benign.
-Hence, in JavaScript, the condition changes that involve option/configuration values deserve attention since they are more probable to be BCs.
+have studied the behavioral breaking change types in Maven projects (Zhang et al., 2022).
+They found that changing execution logic and changing calculation of the output make up over 60% of the total breaking changes collected by them.
+They also studied benign changes that will not cause incompatibilities, such as additional/changed/deleted conditions and branches (over 55% of the total benign changes), additional try-catch statements, and assignment revision.
+However, in our findings about source code patterns of JavaScript behavioral breaking changes, it is very common to change conditions in “if” statements to remove support for some option values or to change evaluation criteria of some results, although it seems benign.
+Hence, in JavaScript, the condition changes that involve option/configuration values deserve attention since they are more probable to be breaking changes.
 4.4. RQ4: Why do Developers Make Breaking Changes in NPM Ecosystem?
-In this section, we summarize six major motivations behind making BCs by carefully analyzing our collected BCs. We show the connection between BC types and motivations in Figure 3.
+In this section, we summarize six major motivations behind making breaking changes by carefully analysis of our collected breaking changes.
+We show the connection between BC types and motivations in Figure 3.
 Figure 3. Distribution of Motivations in Different Types of Breaking Changes
@@
 (469 cases) Some source code is no longer needed, hence developers remove them. There are two cases of this reason:
-  • To unify the approach of an functionality (21 cases).
+  • To unify the approach of a functionality (21 cases).
 The functionality can be achieved by other methods, or by using other packages.
 Therefore, the current implementation can be removed.
-For example, in commit 1354171 of thi-ng/umbrella, developers removed function “writeFile”, and recommend using “writeFile” provided by another package @thi-ng/rstream-log-file since the third-party function can also achieve the same functionality.
+For example, in commit 1354171 of thi-ng/umbrella, developers removed function writeFile, and recommend using writeFile provided by another package @thi-ng/rstream-log-file since the third-party function can also achieve the same functionality.
   • To deprecate and remove an outdated functionality (413 cases).
 Some outdated input (e.g., authentication mechanism) are deprecated thus developers remove their support.
-For example, in commit ad4f1493 of Rest.js, the support for netrc authentication method was removed while others were retained.
+For example, in commit ad4f1493 of octokit/rest.js, the support for netrc authentication method was removed while others were retained.
   • To better organize the project (35 cases).
 Moving part of the functionality out and create a new package will make the project better organized.
-For example, in commit 0fa2830222222https://github.com/jsdoc/jsdoc/commit/0fa2830 of jsdoc/jsdoc, the package @jsdoc/core was reorganized and some methods in @jsdoc/core moved to a new package @jsdoc/cli.
+For example, in commit 0fa2830232323https://github.com/jsdoc/jsdoc/commit/0fa2830 of jsdoc/jsdoc, the package @jsdoc/core was reorganized and some methods in @jsdoc/core moved to a new package @jsdoc/cli.
 Table 4. Typical Rename Operations
@@
 Therefore, after a code change, the identifier names tend to be modified.
-For example, in issue 3403 of angular/angular [^23] https://github.com/angular/angular/issue/3403, the developer complaint that “When I first saw the requestCheck() method on ChangeDetectorRef, I assumed it was how I manually run change detection.
+For example, in issue #3403 of angular/angular [^24] https://github.com/angular/angular/issue/3403, the developer complaint that “When I first saw the requestCheck() method on ChangeDetectorRef, I assumed it was how I manually run change detection.
 Others have assumed this as well”.
@@
 For example, in commit da19583 of adonisjs/adonis-framework [^27] https://github.com/adonisjs/adonis-framework/commit/da19583, developers renamed Context to HttpContext since Context is a commonly used keyword.
-  • To fix typos (2 cases). Simply fixing a typo. For example, renaming “MetricAarmProps” to “MetricAlarmProps”.
+  • To fix typos (2 cases). For example, renaming MetricAarmProps to MetricAlarmProps.
 We also compared the rename operations in JavaScript (and TypeScript) with those in Java since most Java and JavaScript projects adopt the camelCase naming convention.
@@
 In Section 4.3, we conclude that the main reason for renaming identifiers is to make identifier name reflect the actual functionality.
-Therefore, on the one hand, poor identifier name is a noticeable technical debt in JavaScript projects, hence automatic naming can be utilized to provide good identifier names, and on the other hand, in some cases, renaming identifiers is reasonable due to the necessary adaptation to frequent code changes in NPM ecosystem, thus we suggest applying automatic renaming techniques to make identifier names adapt to code evolution.
+Therefore, on the one hand, poor identifier naming is a noticeable technical debt in JavaScript projects, hence automatic naming can be utilized to provide good identifier names, and on the other hand, in some cases, renaming identifiers is reasonable due to the necessary adaptation to frequent code changes in NPM ecosystem, thus we suggest applying automatic renaming techniques to make identifier names adapt to code evolution.
 Several rule-based, static analysis-based, and machine learning-based approaches are proposed to help automatic naming and renaming.
@@
   • They are not code-change-aware.
-Since we have found that many a proportion of Rename BCs are performed to ensure code consistency (Figure 3 in Section 4.3), a renaming technique had better leverage knowledge in prior related code changes to recommend proper names.
+Since we have found that many a proportion of Rename breaking changes are performed to ensure code consistency (Figure 3 in Section 4.3), a renaming technique had better leverage knowledge in prior related code changes to recommend proper names.
 Tools for detecting code with similar functionalities are necessary.
-According to Figure 3, a number of Remove BCs are motivated by approach for unification, i.e., provide a unified API for a specific functionality.
-However, unlike the casualty changes (e.g., removing a class then the code referring to it should also be removed), the relationship between these co-occurring changes is indirect and not simple to detect.
-Besides, in our BC extraction process, we find developers make other small changes for better performance, compatibility, and robustness though they are unrelated to BCs.
+According to Figure 3, a number of Remove breaking changes are motivated by approach for unification, i.e., provide a unified API for a specific functionality.
+However, the relationship between these two or more code changes is indirect and not simple to detect.
+Besides, in our BC extraction process, we find developers make other small changes for better performance, compatibility, and robustness though they are unrelated to breaking changes.
 This also to some degree interferes with BC identification.
@@
   • Identify functionally related code snippets that can be removed together to facilitate clean code.
-  • Distinguish non-breaking changes from breaking changes and casualty changes to provide more convenience for developers.
-Future research can focus on detecting more types of behavioral BCs.
-Syntactic BC detection (e.g., Rename, Remove and Change Signature) is somehow trivial and many refactoring detection-based tools can achieve it.
-For example, APIDiff (Brito et al., 2018), Clirr (Cli, [n.
-d.]), Japicmp (jap, [n.
-d.]) and RevAPI (Rev, [n.
-d.]) can be used for Java projects, PyCompat (Zhang et al., 2020) and Pidiff (pid, [n.
-d.]) can be applied in for Python libraries, and a variant of APIDiff can be used for JavaScript to characterize refactoring history (Brito et al., 2021).
-Hence the future research works should focus on behavioral BCs that cannot be simply identified.
-Existing such approaches for JavaScript are NoRegrets (Mezzetti et al., 2018) and its enhanced version NoRegrets+ (Møller and Torp, 2019).
+  • Distinguish non-breaking changes from breaking changes and ordinary code changes to provide more convenience for developers.
+Future research can focus on detecting more types of behavioral breaking changes.
+Existing approaches for JavaScript behavioral breaking change detection are NoRegrets (Mezzetti et al., 2018) and its enhanced version NoRegrets+ (Møller and Torp, 2019), to the best of our knowledge.
 In Section 1, we have highlighted their reliance on test cases in downstream projects.
-In addition, the two approaches can only detect type-related BCs.
+In addition, the two approaches can only detect type-related breaking changes.
 Specifically, in the first phase, they run client test cases and monitor the flows of values by program instrumentation to build API models (intuitively, type restriction of APIs), and in the second phase, they rerun the client test cases with updated upstream code to check whether type restrictions are changed during execution.
-Therefore, they cannot detect non-type-related BCs, e.g., a BC that changes the semantics of the returned string, while the return type is still string.
-According to Section 4.3, there are five types of behavioral BC actions, where only changing the specifications of return values can directly cause type changes, while other behavioral change may not change object types during execution.
-Therefore, we recommend future works try to deal with various types of behavioral BCs uncovered by this study.
+Therefore, they cannot detect non-type-related breaking changes, e.g., a BC that changes the semantics of the returned string, while the return type is still string.
+According to Section 4.3, there are four types of behavioral BC actions, where only changing the specifications of return values can directly cause type changes, while other behavioral changes may not change object types during execution.
+Therefore, we recommend future works try to deal with various types of behavioral breaking changes uncovered by this study.
 6. Threats to Validity
 Threats to internal validity are mainly related to the clarity of developers’ documentation of breaking changes.
-Few developers might have not explicitly documented the BCs in commit messages, and commit messages (as well as text in issues and pull requests) might not fully reflect developers’ reason for making BCs.
-Also, the discussion and documentation related to BCs may not fully reflect developers’ opinions.
-Second, in our manual analysis, we might still not understand developers’ intentions since many BC commit messages do not clearly explain what and why they make BCs.
-Additionally, we might regard some changes in code diff as BC actions since they are confusing.
-To address this potential threat, we have implemented a double-checking process to verify the accuracy of our analysis results.
-Threats to construct validity are concerned with the errors during breaking change selection.
-While checking the compliance to Conventional Commits standard, we simply use the regular expression, which might incorrectly classify some commit messages.
-Also, in the process of selecting commits for further study (Section 3.3), we might abandon some commits, although they satisfy our selection criteria.
+Few developers might have not explicitly documented the breaking changes in commit messages, and commit messages (as well as text in issues and pull requests) might not fully reflect developers’ reason for making breaking changes.
+Also, the discussion and documentation related to breaking changes may not fully reflect developers’ consideration.
+Second, in our manual analysis, we might still not understand developers’ intentions since many breaking changes’ commit messages do not clearly explain what and why they make breaking changes.
+Additionally, we might regard some changes in code diff as BC-related actions since they are confusing.
+To address this potential threat, we have double-checked the analysis results to ensure the quality.
+Threats to construct validity are concerned with the errors during breaking commit selection.
+In our study, we mainly consider the commits that follow Conventional Commits specification and collect breaking change commits that contain BREAKING CHANGE token, which may miss the commits that not fully follow the specification but are actually breaking.
+To mitigate this, we try our best to collect breaking changes from software documentation, especially changelogs, issues and pull requests (mentioned in Section 3.1).
+Additionally, we might overlook some commits, although they satisfy our selection criteria although we carefully inspect each collected commits (in Section 3.3).
 Threats to external validity refer to the generalizability of our study.
-In our study, while we involve 1,519 BCs extracted from 381 randomly sampled projects from a vast number of popular NPM projects, and these projects cover diverse application fields of JavaScript language, such as utilities, frontend, Web APIs, database, etc., the extracted BCs only account for a small proportion of BCs in the whole NPM ecosystem.
+In our study, while we involve 1,519 breaking changes extracted from 381 randomly sampled projects from a vast number of popular NPM projects, and these projects cover diverse application fields of JavaScript language, such as utilities, frontend, Web APIs, database, etc., the extracted breaking changes only account for a small proportion of breaking changes in the whole NPM ecosystem.
 Besides, in this study we do not consider less popular projects since they have short development history and it is less probable to yield some patterns from these projects.
-However, the BC practices in less popular projects may still differ from those in popular projects since they are not so well maintained.
+However, the BC practices in less popular projects may still differ from those in popular projects.
+As for another threat, the Reid et al.’s dataset (Reid, 2020) used in our study was released in 2020 (four years ago), and might not include popular JavaScript and TypeScript released after 2020.
+We manually checked 500 NPM libraries with the largest number of dependents and found that between them [^29] We use the list in https://leodog896.github.io/npm-rank/PACKAGES.html, which is updated automatically (We accessed the page in September 2024).
+only 16 were initially published after 2020.
+Hence Reid et al.’s dataset probably covers most popular JavaScript and TypeScript projects, and this threat might not affect the generalizability too much.
 Threats to conclusion validity relate to insufficient high-quality BC commits to support our findings.
-To mitigate the threats, we try our best to extract breaking changes from commit messages, issues, pull requests and changelogs to construct our BC dataset.
-We also carefully inspect our collected BCs and retain those related to JavaScript source code changes and explicitly documented by developers, which can help increase the quality of our dataset.
+To mitigate the threats, we try our best to extract breaking changes from commit messages, issues, pull requests and changelogs to construct our breaking change dataset.
+We also carefully inspect our collected breaking changes and retain those related to JavaScript source code changes and explicitly documented by developers, which can help increase the quality of our dataset.
 7. Related Work
@@
 7.3. Breaking Change Detection and Analysis on Other Platforms
-Besides JavaScript, a lot of studies focus on detecting and analyzing BCs for other programming languages, especially Java and Python, and their techniques may be learned by BC detection for JavaScript.
+Besides JavaScript, a lot of studies focus on detecting and analyzing breaking changes for other programming languages, especially Java and Python, and their techniques may be learned by breaking change detection for JavaScript.
 Brito et al.
@@
 Regarding non-syntactic breaking changes, to the best of our knowledge, Zhang et al.
-proposed Sembid (Zhang et al., 2022) to detect behavioral BCs by measuring the semantic difference of call graphs between old and new programs: if a code change’s semantic difference is larger than a threshold and not identified as benign change, it is classified as a BC.
-However, their used semantic diffs reflect the structural changes of programs, while many semantic BCs do not need many of those changes, e.g., only changing one condition, and adding a small change to the output string.
-Additionally, as the authors have pointed out, Sembid cannot detect high-level BCs such as re-implementation of a method.
+proposed Sembid (Zhang et al., 2022) to detect behavioral breaking changes by measuring the semantic difference of call graphs between old and new programs: if a code change’s semantic difference is larger than a threshold and not identified as benign change, it is classified as a BC.
+However, their used semantic diffs reflect the structural changes of programs, while many semantic breaking changes do not need many of those changes, e.g., only changing one condition, and adding a small change to the output string.
+Additionally, as the authors have pointed out, Sembid cannot distinguish semantic breaking changes from non-breaking changes such as re-implementation of a method.
 Therefore, directly adapting Sembid to JavaScript projects is not suitable.
-Additionally, with the help of BC detection techniques, many works analyzed the impacts of breaking changes to downstream projects.
+Additionally, with the help of breaking change detection techniques, many works analyzed the impacts of breaking changes to downstream projects.
 For example, Jayasuriya et al.
-(Jayasuriya et al., 2023, 2024) investigated the impacts of breaking changes (especially behavioral breaking changes) to client applications in Maven ecosystem.
+(Jayasuriya et al., 2023, 2024) investigated the impacts of breaking changes (especially behavioral breaking changes) to client applications in the Maven ecosystem.
 They concluded with many findings, e.g., 11.58% of the dependency updates contain breaking changes that impact the clients and 2.30% version updates have behavioral breaking changes that impacted client tests.
@@
 8. Conclusion and Future Work
-In this study, for better comprehension of breaking changes in the NPM ecosystem and enhancing BC-related tools, we conduct an empirical study to bridge the knowledge gap from three aspects, with our carefully constructed breaking change dataset (1,519 BCs in total) sampled from a large set of popular NPM projects.
-We find that 97% of BCs detected by regression testing can be covered by developers’ documentation, which proves that extracting BCs from documentation is reasonable.
-We then summarize the JavaScript and TypeScript specific syntactic BCs and how developers make behavioral BCs, yielding many findings.
-Besides, we also investigate the reasons behind BCs in JavaScript and conclude with a taxonomy, which extend the previous works on motivations behind breaking changes.
-Based on our empirical findings, we provide actionable implications for future research, e.g., applying automatic renaming and naming techniques in JavaScript projects, and detecting code with similar functionalities, etc.
+In this study, for better comprehension of breaking changes in the NPM ecosystem and enhancing breaking change-related tools, we conducted an empirical study to bridge the knowledge gap from three aspects, with our carefully constructed breaking change dataset (1,519 breaking changes in total) sampled from a large set of popular NPM projects.
+We found that 95.4% of the breaking changes detected by regression testing can be covered by developers’ documentation, which proves that extracting breaking changes from documentation is reasonable.
+We then summarized the breaking changes in the NPM projects that are specific to JavaScript and TypeScript, and how developers make behavioral breaking changes, which yield many findings.
+Besides, we also investigated the reasons behind breaking changes in JavaScript and conclude with a taxonomy, which extend the previous works on motivations behind breaking changes.
+Based on our empirical findings, we provided actionable implications for future research, e.g., applying automatic renaming and naming techniques in JavaScript projects, and detecting code with similar functionalities, etc.
 In the future, we want to collect additional sources such as online discussions, and employ alternative approaches like dynamic analysis to gain a deeper understanding of breaking changes in the NPM ecosystem.
-We plan to investigate how to automatically identify the breaking changes in NPM projects.
-We consider combining static analysis and dynamic analysis techniques to enhance the existing BC detection approaches.
-We also consider improving BC detection with the help of Large Language Models since they show great performance improvements in many software engineering related tasks and can utilize the semantics in source code.
+We plan to investigate how to automatically identify the breaking changes in the NPM projects.
+We consider combining static analysis and dynamic analysis techniques to enhance the existing breaking change detection approaches.
+We also consider improving breaking change detection with the help of large language models since they show great performance improvements in many software engineering related tasks and can utilize the semantics in source code.
+Acknowledgements.
+This research/project is supported by the National Science Foundation of China (No.62372398, No.72342025, and U20A20173), the Fundamental Research Funds for the Central Universities (No.
+226-2022-00064), and the National Research Foundation, under its Investigatorship Grant (NRF-NRFI08-2022-0002).
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore.